modification dates for phil's home page:
- 06may13
added page for millstone hill wideband feed.
- 06may13
System performance for the apr13
- 06may13 fractional
occurrence of rfi for apr13
- 06may13 Alfa
Tsys for the month of apr13
- 06may13
ages (a2048) monitoring of data for month of apr13
- 06may13
sitepower monitoring for month of apr13
- 06may13 tiedown
tensions for apr13
- 06may13 azimuth
motor torques vs velocity for apr13
- 02may13
tiedown tensions for apr13. When do the tiedowns lose tension..
- 29apr13
measuring the punta Borinquen radar blanking in the AO
direction.
- 29apr13
check lbw filterbank on 22apr13 by switching between filters.
- 24apr13
new punta borinquen carsr 1274,1332 Mhz radar
- 23apr13
added lbw rfi filter bank page
- 22apr13:
a2765 dps processing on 21apr13 using xband.jumps in adjacent
bands.
- 13apr13
System performance for the mar13
- 13apr13 fractional
occurrence of rfi for mar13
- 13apr13 Alfa
Tsys for the month of mar13
- 13apr13
ages (a2048) monitoring of data for month of mar13
- 13apr13 tiedown
tensions for mar13
- 13apr13 azimuth
motor torques vs velocity for mar13
- 11apr13:
cband 4920 rfi present 06apr13-09apr13.
- 10apr13
a2749. started daily monitoring of their data. polB had drift at
start of run.
- 09apr13:
a2749 using lbw had polB drift by 2% for the first 45
minutes of run.
- 06apr13
cband 5-6 ghz azswings done 28aug12.
- 04apr13
added alfa shutter page.
- 03apr13
started ao19 phased array feed page.
- 03apr13
netgear gsm7238fs network switch. Measured 17oct12 by luis in
the screen room.
- 03apr13
dome 430 polA tsys close to 400 K during afternoon Tsys
measurements.
- 01apr13
netgear prosafe gs110tp network switch. Emissions Measured
06sep2 by luis in the screen room
- 01apr13
netgear prosafe gs110t network switch. Emissions measured16oct12
by luis in the screen room
- 31mar13
Softek ivuloto device at visitors center. measured 15oct12 by
luis in screen room.
- 28mar13
p2787 pulsar observing (27mar12) had spectral jumps from digital
filter in mocks.
- 27mar13
rserv2 bulldozer fftw times using kernel 3.4.33. multithreaded
ffts improved.
- 27mar13
6900-7050 rfi seen 11mar13,16mar13 by chris salter
- 25mar13
xmitting tone from control room patio: Mapping screen room
birdie strength to detectability on the telescope
- 23mar13
azimuth peak with za=17deg and transmitting from control room
patio.
- 23mar13
elo touch screen cash register emissions: in screen room and on
telescope.
- 21mar13
added daily monitoring of site power
- 12mar13
rfihilltop will plot multiple days of rfi hilltop monitor data.
dynamic spectra or daily averages.
- 12mar13
System performance for the feb13
- 12mar13 fractional
occurrence of rfi for feb13
- 12mar13 Alfa
Tsys for the month of feb13
- 12mar13
ages (a2048) monitoring of data for month of feb13
- 12mar13 tiedown
tensions for feb13
- 12mar13 azimuth
motor torques vs velocity for feb13
- 08mar13
added page with radar calibration satellites.
- 06mar12
measure rfi from touch screen cash register in screen room (for
cafeteria).
- 05mar12
trelectronics LB65 distomat tested for rfi in screen room.
- 04mar13 td
4 position jumps to 24 inches and back on 02,03mar13
- 28feb13
look at new rf monitor installed at the visitors center.
- 23feb12
galfacts see wideband rfi in alfa on 14feb13.
- 21feb12
updated software measurements page. fftwV3.3.3 and added
some comparisons.
- 15feb13
added amd software documentation page.
- 15feb13
rserv2 (amd 6272 64 core machine) fftw timing results using avx
instructions.
- 14feb13
sine wave test 04feb13 with aeronomy rdev device.
- 13feb13
cband calibration run after dewar reinstalled in dome.
- 10feb13
added t2686 (herb and frank) userproj page. plasma line images
once an hour for run.
- 07feb13
System performance for the jan13
- 07feb13 fractional
occurrence of rfi for jan13
- 07feb13 Alfa
Tsys for the month of jan13
- 07feb13
ages (a2048) monitoring of data for month of jan13
- 07feb13 tiedown
tensions for jan13
- 07feb13 azimuth
motor torques vs velocity for jan13
- 04feb13 added
daily plots of dewar bias voltages and currents for single pixel
receivers.
- 02feb13:
cband polB instability at crosshead freq. also seen in stage 2
drain current.
- 01feb13:
x111 dynamic spectra for lband nov/dec12
- 31jan13
added manuals for "new" site power substation controller.
- 31jan13
added schematics, manuals for building II A/C unit.
- 31jan13
building II A/C schematics
- 28jan13:
spectrum analyzer at building 2 a/c unit shows rfi.
- 28jan13
preliminary 327 Mhz results for building 2 shutdown on 24jan13.
- 24jan13
azimuth encoder difference before/after az encoder 2 changed on
17jan13.
- 24jan13
mock spectrometer b4sXg0 PI FIFO overflow during galfacts
observing.
- 23jan13:
az position for peak signal from building 2.
- 14jan13
simulate puppi a/d sampling errors.
- 13jan13
calibration data jun12-dec12. updated pointing offsets.
- 11jan13
nov12 horn position survey.
- 10jan13
yearly platform wind speed plots. added 2012
- 10jan13
yearly platform temperature plots. added 2012.
- 10jan13
System performance for the dec12
- 10jan13 fractional
occurrence of rfi for dec12
- 10jan13 Alfa
Tsys for the month of dec12
- 10jan13
ages (a2048) monitoring of data for month of dec12
- 10jan13 tiedown
tensions for dec12
- 10jan13 azimuth
motor torques vs velocity for dec12
2012 (top)
- 13dec12
performance of 430 dome cavity filter (with notches) in front of
dewar.
- 13dec12
System performance for the nov12
- 13dec12 fractional
occurrence of rfi for nov12
- 13dec12 Alfa
Tsys for the month of nov12
- 13dec12
ages (a2048) monitoring of data for month of nov12
- 13dec12 tiedown
tensions for nov12
- 13dec12 azimuth
motor torques vs velocity for nov12
- 12dec12
oct-nov12 lbw Tsys increase. calibration data shows it's not a
problem with the cals.
- 10dec12.
sbn compare old, new (out of focus) , new (in focus) calibration
runs.
- 10dec12
07dec12 lbw cooldown. 16k stage took 1 hour.
- 07dec12
a2704 430 dome continuum mapping show polA ripples (after filter
changed).
- 06dec12
430gr azswings on 30nov12. polA filter removed, tv
station intermods, az dependence, time periodic birdies
- 04dec12
430gr, 19nov12 take spectra while moving radar,radio astronomy
switch. Ripple in both positions.
- 03dec12
lbw dewar 16k stage took < 36 hours to cool down
- 27nov12
System performance for the oct12
- 27nov12 fractional
occurrence of rfi for oct12
- 27nov12 Alfa
Tsys for the month of oct12
- 27nov12
alfalfa (a2010) monitoring of data for month of oct12
- 27nov12
ages (a2048) monitoring of data for month of oct12
- 27nov12
alfa stability (using a2010 date) PolA/polB ratios for oct12
- 27nov12 tiedown
tensions for oct12
- 27nov12 azimuth
motor torques vs velocity for oct12
- 26nov12
430 dome x111 data from oct,nov12 show the polA ripple coming
and going.
- 23nov12
new sbn receiver. calibration runs comparing old and new
receiver.
- 22nov12 Pr Law title
23, Part I chapter 11: Radio-Astronomy Facility zoning act.
- 22nov12 fcc Arecibo
coordination zone document
- 21nov12
cal values for new sbn receiver measured and installed.
- 16nov12
added rabbit bl2100 sbc documentation (used in rf radiation
monitor).
- 14nov12
polA ripple in lbw from 1370 highpass filter rf switches.
- 14nov12
Tsys contribution various configurations of sbn postamps.
- 09nov12
added hardware page for cable car.
- 06nov12 a2713 dps
with cband. 1 mock with glitches. 1 with ripple for 1 pattern.
- 02nov12
preliminary results from nov12 horn survey
- 31oct12
cband polA stage 2 drain current has large rms with amp#18, but
not amp#19
- 24oct12
hilltop rfi monitor. Added schematics and program documentation.
- 22oct12
430 dome polB jennings relay causing little power from telescope
to get to amps.
- 22oct12
a2415 430 mapping
- 15oct12
430 dome polB gain is 10% of the expected value.
- 11oct12
System performance for the sep12
- 11oct12 fractional
occurrence of rfi for sep12
- 11oct12 Alfa
Tsys for the month of sep12
- 11oct12
alfalfa (a2010) monitoring of data for month of sep12
- 11oct12
ages (a2048) monitoring of data for month of sep12
- 11oct12
alfa stability (using a2010 date) PolA/polB ratios for sep12ii
- 11oct12 tiedown
tensions for sep12
- 11oct12 azimuth
motor torques vs velocity for sep12
- 10oct12
alfalfa sep12 data shows long recovery times in beam1b with
radars.
- 09oct12
mock spectrometer test with cband and pshift=0x1555 gets rid of
spectral jumps.
- 09oct12
added page describing mock spectral jumps spaced by 2^n channels
- 07oct12
mock spectrometer fgpa dma overflow on b0s1g0 during a2130
- 04oct12
a2708 freqsw partial results.
- 02oct12
p2679 bits flips in mock highres mode.
- 01oct12
added table of all pdev spectrometers and their status.
- 28sep12
azwings with cband hi 6-7 Ghz shows rfi az dependence
- 28sep12 a2708
freqsw partial results.
- 27sep12 cband
SEFD ratio PolA/Polb after polA amp change shows Tsys is
different
- 27sep12 a2708
freqsw partial results.
- 20sep12 430gr
TsysPolA is jumping by 10 to 15 K.
- 19sep12
installed new cband high cal values that were measured back on
28jun12.
- 17sep12
installed new cband cal values that were measured back on
28jun12.
- 14sep12
09sep12-11sep12: 2nd lo synthesizers 1-4 are stuck at at a fixed
frequency.
- 11sep12
puppi: debugging 32 chan, 200 Mhz raw mode
- 05sep12
atm power profiles show cal dropouts.
- 03sep12
System performance for aug12
- 03sep12 fractional
occurrence of rfi for aug12
- 03sep12 Alfa
Tsys for the month of aug12
- 03sep12
ages (a2048) monitoring of data for month of aug12
- 03sep12 tiedown
tensions for aug12
- 03sep12 azimuth
motor torques vs velocity for aug12
- 30aug12
cband cable from bias chassis to dewar replaced on 29aug12.
- 29aug12 5275
Mhz Expernet (xairnet) internet test from jonathan
friedman's house.
- 28aug12
27aug12 sbh bias box moved to cband. polA stage 2 drain current
problem remains.
- 28aug12 added
page with 2012 cband polA instabilities.
- 28aug12
interim correlator sbc4A band had a larger tsys than other sbc.
Function of bandwidth.
- 24aug12 a2471 brown
dwarf bursts at cband (alex).. processed data 2011,2012.
- 22aug12 monitor
wind speeds during storm isaac
- 20aug12
cband the rms the rf power for polA increases while stage 2 bias
drain current jumps around.
- 16aug12
xband cal values remeasured on 01jul12 installed 16aug12.
- 16aug12 added xband
stokes plots
- 12aug12
lband wide cal values remeasured on 28jun12. installed 12aug12.
- 08aug12
sband tx. plot old and new voltage to power lookup tables for
antenna power.
- 08aug12
added page " measuring the sband transmitted power."
- 08aug12 fractional
occurrence of rfi for jul12
- 08aug12 Alfa
Tsys for the month of jul12
- 08aug12
ages (a2048) monitoring of data for month of jul12
- 08aug12 tiedown
tensions for jul12
- 08aug12 azimuth
motor torques vs velocity for jul12
- 31jul12
sband transmitter. plot lro power meter values vs "antenna
power" from 723 program.
- 31jul12
cband polB bias values jul12 show stage 2 currents polB jumping.
- 05jul12 fractional
occurrence of rfi for jun12
- 05jul12 Alfa
Tsys for the month of jun12
- 05jul12
ages (a2048) monitoring of data for month of jun12
- 05jul12 tiedown
tensions for jun12
- 05jul12 azimuth
motor torques vs velocity for jun12
- 05jul12
added rfi page for yiye avila's tower on roncador.
- 02jul12
check cband receiver (31may12) after it was reinstalled in the
dome.
- 01jul12 bench
marked intel avx routines using fftw-3.3.2 on megs3
- 01jul12
glitches in rf from alfa monitoring are only present with bias
monitoring.
- 29jun12
added plots of rf monitor at visitor center: feb09-jun12
- 29jun12 refit dut1
(ut1-utc) for leap second second 01jul12
- 26jun12
sband narrow stokes calibration results (01jan07 -> 31may12)
- 26jun12
sband wide stokes calibration results (may03->31may12)
- 25jun12
cband stokes calibration results (16nov10 thru 31may12)
- 25jun12 sband
high stokes calibration results (28jun04 thru 31may12)
- 25jun12
lband wide stokes calibration results (16jan11 thru 27may12).
- 25jun12 cband
high stokes calibration results.
- 20jun12
pointing error as we tilt the platform. Computed and measured
values.
- 19jun12
12 meter first light.
- 18jun12
pointing error as we move the platform vertically. computed and
measured offsets.
- 18jun12
rfi in screen room from 12meter camera (measured by luis)
- 16jun12
tiltsensor az swings and za strips
- 16jun12
tiltsensor az swings
- 12jun12
p2030 data show dewarmonitoring and .838Hz birdie (but no 1.5 hz
rfi)
- 09jun12
added page tracing sband body current signal through the
schematics.
- 08jun12
distomat 6 loses alignment when test distomat mounted on it.
- 08jun12
added page describing cima changes for turning on,off alfa
bias monitoring.
- 05jun12
model18 verify completed, model18 installed.
- 04jun12 fractional
occurrence of rfi for may12
- 04jun12 Alfa
Tsys for the month of may12
- 04jun12
ages (a2048) monitoring of data for month of may12
- 04jun12 tiedown
tensions for may12
- 04jun12 azimuth
motor torques vs velocity for may12
- 04jun12 fractional
occurrence of rfi for apr12
- 04jun12 tiedown
tensions for apr12
- 04jun12 azimuth
motor torques vs velocity for apr12
- 02jun12
alfa dewar monitoring is creating periodic gain glitches in the
receiver.
- 28may12 a2130
(galfacts) see galileo satellite interference at 1278.7 Mhz.
- 24may12
tilt sensor az swings. cband receiver in the lab
- 23may12:
strobe lights coming from yiye avila's tower on mount roncador.
- 22may12
added plots for a2430 browndwarf stokesV bursts (from dec09)
- 19may12 td8
releases by itself.
- 16may12 r2696 (
lro) run on the moon.
- 14may12 r2696 lro
test run showing chirped transmitter pulse.
- 12may12 cband
stage 2 polA drain current, still jumping around after switch to
stage 3 circuitry.
- 11may12 faa radar
ipps change by .3%. Dip switches for timing probably set
incorrectly.
- 11may12 cband
crosshead oscillation causes a 1% gain jump in Tsys.
- 10may12
cband stage 2 polA drain current jumps after bias adjusted on
09may12.
- 05may12
2394.5-2403.5 Mhz rfi coming from yiye avila's tower (roncador).
- 05may12
3rd harmonics of fm stations 102.9.104.1 from route 10.
- 03may12
calibration scans using alfa after reinstallation on 20apr12.
- 30apr12
tilt sensor measurements after td12 jack replaced.
- 24apr12 ao users
committee talk for 2012.
- 23apr12 td12 jack
replaced. New limits set
- 20apr12 327
rcvr cavity filter (before dewar) bandpass shape.
- 20apr12
added tiedown grounding page. Include resistance to ground
measurements for td12
- 19apr12
td12 stops with foldback error. Needs large torques to move..
jack to be replaced.
- 18apr12
a2335 lbw 14apr12. rain causes lbw tsys to increase.
- 10apr12 fractional
occurrence of rfi for mar12
- 10apr12 tiedown
tensions for mar12
- 10apr12 azimuth
motor torques vs velocity for mar12
- 04apr12
simulate linefeed gain vs za. Compare with measured values.
- 03apr12 can we
replace the 430tx with a ring of amiser panels at the edge
of the dish?
- 31mar12
terminate input to 430 rcvr at various places before
amp looking at polA ripple.
- 28mar12
Plots showing ch430 transmitter modAnode not staying
completely off after beam off.
- 28mar12
430 dome polA ripple while receiver on load.
- 22mar12 cband
microwave interference at 5990-6020.
- 20mar12
added kollmorgen motor repair log.
- 17mar12
polA ripple 430 dome continues.
- 14mar12 fractional
occurrence of rfi for feb12
- 14mar12 Alfa
Tsys for the month of feb12
- 14mar12
ages (a2048) monitoring of data for month of feb12
- 14mar12 tiedown
tensions for feb12
- 14mar12 azimuth
motor torques vs velocity for feb12
- 12mar12
scramPuppi. monitor scramNet multicast packets and load
puppi shared memory.
- 08mar12 2140-2150 rfi
from Advanced Wireless System (AWS).
- 05mar12 pointing
model18 created. Not yet installed.
- 29feb12 tiedown daily
monitoring now includes status bits for all 3 tiedowns.
- 28feb12 refit ut1 -
utc polynoimial for mar12 thru jun12
- 26feb12 tiedown 1 sec
tick fiber distribution switched to 3 separate fibers.
- 24feb12 fractional
occurrence of rfi for jan12
- 24feb12 Alfa
Tsys for the month of jan12
- 24feb12
ages (a2048) monitoring of data for month of jan12
- 24feb12 tiedown
tensions for jan12
- 24feb12 azimuth
motor torques vs velocity for jan12
- 22feb12
mock spectrometer dump status timeouts resolved.
- 20feb12
tiedown 4 tmscounter time drifting.
- 20feb
12 b6sNg0 (pdevs7) continues to timeout on while taking
data.
- 17feb12
puppi first light 14feb12 with lbw. Radars are saturating.
- 16feb12
azswings with tilt sensors to check "new" td positions.
15,16may12.
- 13feb12
10feb12 changed td a/d signal conditioning resistor for spdmon
from 46.2k to 22k.
- 11feb12
debugging problems with pdev: integration stops early
- 08feb12
added schematics for jon hagens 430 transmitter manual
- 07feb12
hp87130 errors rfattn polB does not switch.
- 02jan12
move td's, read output voltages, and then compare with readback
values.
- 01feb12
updated scramnet and multicast data documentation.
- 29jan12
.shs clp processing times using 45 threads on rserv1.
- 29jan12
16jan12
reflection coef measurements of 430 Tx waveguide. Reflections
and losses.
- 28jan12
table
of distance along 430 waveguide using waveguide
section lengths from hagen's manual
- 27jan12 steps
taken when making a pointing model.
- 27jan12
azencoder
1 replaced. pictures of how far the pointer at 270 moved.
- 25jan12
tiedown
commanded and amplifier readback speed mismatch.
- 23jan12
azimuth
encoder difference after enc1 coupling tightened.
- 20jan12
corcom
ac power line filters.
- 10jan12
yearly
platform temperature plots for 2011
- 10jan12
yearly
windmeter meter plots for 2011
- 10jan12 fractional
occurrence of rfi for dec11
- 10jan12 Alfa
Tsys
for the month of dec11
- 10jan12
ages (a2048) monitoring of data for month of dec11
- 10jan12 tiedown
tensions
for dec11
- 10jan12 azimuth
motor
torques vs velocity for dec11
2011 (top)
- 13dec11
description
of merging multiple sps atm ipps for T2574
- 11dec11
10dec11
sband log plots
- 07dec11
alfa
filter
bank switching test.
- 06dec11
replacing the tiedown control system.
- 05dec11
rediscovering
the Galaxy using atm az swings.
- 05dec11
rfi
from
vlfant from sri. azswing (narrow band) around 1400Mhz.
- 05dec11 fractional
occurrence of rfi for nov11
- 05dec11 Alfa
Tsys
for
the month of nov11
- 05dec11
alfalfa
(a2010) monitoring of data for month of nov11
- 05dec11
ages (a2048) monitoring of data for month of nov11
- 05dec11
alfa stability (using a2010 date) PolA/polB ratios for nov11ii
- 05dec11 tiedown
tensions
for
nov11
- 05dec11 azimuth
motor
torques
vs velocity for nov11
- 30nov11 fractional
occurrence of rfi for oct11
- 30nov11 Alfa
Tsys
for
the month of oct11
- 30nov11
alfalfa
(a2010) monitoring of data for month of oct11
- 30nov11
alfa stability (using a2010 date) PolA/polB ratios for oct11ii
- 30nov11 tiedown
tensions
for
oct11
- 30nov11 azimuth
motor
torques
vs velocity for oct11
- 28nov11
alfa
beam
2a
shows some jumps in ages data.
- 27nov11
added
page with plots of sband transmitter log files.
- 25nov11
azimuth
encoder
difference
after az encoder 2 replaced.
- 25nov11
added
daily
summary
plot for az,gr,ch.
- 19nov11
sband
antenna
power
vs lro power meter.
- 15nov11 page
for programmable ac power strip:
- 08nov11
possible 2 mhz comb in 327 receiver.
- 08nov11
rfi from alfa motor controller at 327 Mhz measured
03nov11.
- 07nov11
Looking
for
rfi
from the visitor center by shutting off its power (01nov11)
- 07nov11
xmitting
from
the
visitor center platform to find the azimuth peak (01nov11)
- 04nov11
puntas
salinas
creates
intermods in a2130 data 02nov11.. added a2130 overflow counts
- 31oct11
added documentation for corstripch interim cor stripchart
program.
- 30oct11
td
4
makes
load noises 27,29oct11.
- 30oct11
i/o
rates
29nov11
on pdevs discs during p2627 (VoltSampling),p2010,a2130(galfacts)
- 27oct11
fit
C0,1az,
3az
pointing errors for alfa after reinstallation in sep11.
- 26oct11
a2621 on/off and dps patterns using 800 Mhz rcvr.
- 24oct11
alfa
calibration
scans
sep_oct11 after alfa reinstalled.
- 22oct11
19oct11
cleaning
alfa
fiber connectors jumped bm1a Tsys to 150K.
- 20oct11
174
khz
ripple
in alfa fibers after fiber connectors were cleaned (data from
19oct11)
- 18oct11
174
khz
ripple
in alfa fibers measured using galfa data from 14oct11
- 18oct11 added page for
galfa spectrometer software
- 17oct11
2401.5 Mhz rfi moves down to 2396.5 on 08oct11.
- 17oct11 fractional
occurrence of rfi for sep11
- 17oct11 Alfa
Tsys
for
the
month of sep11
- 17oct11
alfalfa
(a2010) monitoring of data for month of sep11
- 17oct11
alfa stability (using a2010 date) PolA/polB ratios for sep11ii
- 17oct11 tiedown
tensions for sep11
- 17oct11 azimuth
motor
torques
vs
velocity for sep11
- 14oct11
added
a2130
caldeflection
plots for start of oct11 run (07-13)
- 12oct11
added monthly plots of ism band from hilltop monitoring. Show
2395 rfi.
- 11oct11
Beam
4a
tsys
while
dana took data with the spectrum analyzer.
- 08oct11
added
rfi
page
for
AC units above buildings I and II.
- 07oct11
alfa
a2010
total
power:
bm4a with low drain current and bm6b failure
- 04oct11
alfa
bm6b
power
drops
out 03oct11.
- 02oct11
Aeronomy
data
page
with
clp clutter removal.
- 30sep11
6595 rfi, 30 Mhz wide seen by a2622 using cband Hi.
- 29sep11
1408
Mhz
rfi
coming
from AC unit on roof of building I.
- 27sep11
alfa
bm5a
dropout
seen
by alfalfa.
- 26sep11
1408
Mhz
rfi
in
alfa data.
- 22sep11 moved
software documentation on home page to software.html.
reorganized it.
- 21sep11
added ryan seals thesis about the echotek card and digital
receiver.
- 14sep11
mock
time
domain
sampling
on 1713+0747 at lband taken 11sep11
- 11sep11 8
bit rms vs adjpower value for time domain sampling.
- 10sep11 fractional
occurrence of rfi for aug11
- 10sep11 tiedown
tensions for aug11
- 10sep11 azimuth
motor
torques
vs
velocity
for aug11
- 08sep11
sbn,alfa calibration runs to check alfa pointing after
reinstallation.
- 01sep11 added
alfa temps during cool down sequence on 16mar11
- 24aug11
430
power meter data (ladybug)
- 24aug11 wind
speeds during tropical storm irene.
- 18aug11
table of intel IPP lib FFT times on aserv11 (my benchmark)
- 17aug11 lbw
1360.015 birdie. azswing shows not in the dome.
- 16aug11
430 dome polA ripple remains when radar/radioastronomy switch
bypassed (11aug11)
- 16aug11 lbw
1360 narrow band rfi.
- 15aug11
.pdev
file
documentation
for
mock
spectrometer.
- 15aug11
interim
cor
sbc1
polB
spike
in center of band for bw <=12.5Mhz
- 15aug11
intel performance primitives library fft times on aserv11 (intel
benchmark)
- 10aug11 fractional
occurrence of rfi for jul11
- 10aug11 tiedown
tensions for jul11
- 10aug11 azimuth
motor
torques
vs
velocity
for
jul11
- 10aug11
short
duration broadband rfi seen by p2558 in 327 rcvr.
- 06aug11
rdev data on telescope shows overflow wraparound.
- 05aug11
testing
tamarra's
rdev
device
using
a
sine
wave
- 02aug11
fftw timing benchmark for aserv11 (nehalem) dual quad core
machine.
- 01aug11 430dome
polA ripples in a2599 dps data.
- 01aug11
added
diagram
(list)
of
pdev
control
programs
and
what they do).
- 31jul11
started online documentation for .pdev file formatt for mock
spectrometer
- 28jul11 started
online documentation for mock spectrometer psrfits file format.
- 28jul11 fractional
occurrence of rfi for jun11
- 28jul11 tiedown
tensions for jun11
- 28jul11 azimuth
motor torques vs velocity for jun11
- 01jul11
galfacts beam2a caldeflection jumps by 4% 28,29jun11
- 01jul11
pointing corrections after removing 1az term from platform.
- 30jun11
added hagens 430 tx manual.
- 28jun11
tilt
sensor
2
azswings
22jun11
to
check
1
az term.
- 26jun11
galfacts
(a2130)
adc
and
pfb
overflows
for
jun11
vs za.
- 25jun11
TRelectronics
distomat
(llb500)
rfi
test
in
screen
room
15,23jun11
- 23jun11
mock spectrometer. tables of minimum allowable integration
times.
- 23jun11
430gr polA ripples return.
- 22jun11
sbc3,4
of
interim
correlator
Tsys
increases
by
3K.
halfband 3,4 +5volt supply error.
- 17jun11 fit
1az term to new tiedown positions and add to model
- 15jun11 fractional
occurrence of rfi for may11
- 15jun11 tiedown
tensions for may11
- 15jun11 azimuth
motor torques vs velocity for may11
- 10jun11
tested compressor monitor on telescope 06jun11 using 327
receiver
- 06jun11
tiedown
8 failure 110602. fixed 110606.
- 05jun11
a2335
(on 03jun11) sees lbw polB radar compression with long recovery
time.
- 05jun11
plotted
total power vs time for a2335 data mar11 thru 03jun11
- 25may11
a2335
see 1367 Mhz rfi on 18may11 (punta salinas chan 18)
- 24may11
added
anritsu spectrum analyzer manuals.
- 23may11
adjusting stage 3 drain voltage for alfa bm1
fixed bm1a jumps
- 20may11
a2130 alfa beam1b caldeflection jumps with 20k stage diode temp.
- 18may11
rfi test accurange3000 distomat using 327 rcvr. (data taken
20apr11)
- 17may11
Tightening the polA Andrews connector fixed dome 430 polA ripple
- 16may11
Most
active requestors of naic http daemon jan11-may11
- 13may11
compute new td offsets with tilt sensor and az swings.
- 12may11 tilt
sensor az swings to compute new td positions to level platform.
- 11may11 fractional
occurrence of rfi for apr11
- 11may11 Alfa
Tsys for the month of apr11
- 11may11
alfalfa
(a2010) monitoring of data for month of apr11
- 11may11
ages (a2048) monitoring of data for month of apr11
11may11
alfa
stability
(using
a2010
date)
PolA/polB
ratios
for
apr11
-->
- 11may11 tiedown
tensions for apr11
- 11may11 azimuth
motor torques vs velocity for apr11
- 11may11
testing trelectronics distomat
- 11may11
rfi check during 1999 site power shutdown.. 430 and lband data.
- 09may11
testing accurange 3000 distomat replacement 28apr11.
- 05may11
a2130 caldeflection. added 13,14,15 apr11.
- 05may11
small jerks in turret motion while rotating.
- 05may11 430
dome polA shows irregularities in its bandpass.
- 04may11
a2130 caldeflection vs za. beam 4b increases at high za.
- 02may11
a2130 galfacts caldeflection shows alfa gain stability.
- 01may11
galfa bandpass ripples caused by fiber reflections.
- 26apr11
testing averaging of accurange 3000 distomat replacement.
- 26apr11
instabilities in alfa beam 1b seen by a2010,a2130,a2048.
- 19apr11 refit
polynomial for dut1 (utc-ut1) apr11 thru apr12
- 15apr11
explain
why
mock
single
pix
does
not
support
winking
cals
tied
to mock buffers.
- 15apr11
rfi testing of compressor monitor system in the screen room.
- 08apr11
rfi testing of iLon smartserver 2000 AC controller for
buildings 1,2
- 07apr11
rfi testing of accurange3000 distomat.
- 07apr11 added
stokes calibration 800 rcvr using 3C138 on 11mar11.
- 06apr11
az dependence of 327 rfi (fm stations) data taken 09mar11.
- 06apr11 fractional
occurrence of rfi for mar11
- 06apr11 Alfa
Tsys for the month of mar11
- 06apr11
alfalfa
(a2010) monitoring of data for month of mar11
- 06apr11
ages (a2048) monitoring of data for month of mar11
- 06apr11
alfa stability (using a2010 date) PolA/polB ratios for mar11
- 06apr11 tiedown
tensions for mar11
- 06apr11 azimuth
motor torques vs velocity for mar11
- 04apr11
a2471 cband bursts. Look at stokes I variation various days.
- 31mar11
radar intermods in galfa spectrometer.
- 23mar11
mock time domain sampling of 1937+21 at lbw, and 430Mhz.
- 22mar11
a2335..
total power plots for some of the mar11 data.
- 15mar11
1357.2
Mhz rfi in lbw.
- 15mar11 fractional
occurrence of rfi for feb11
- 15mar11 Alfa
Tsys for the month of feb11
- 15mar11
alfalfa
(a2010) monitoring of data for month of feb11
- 15mar11
ages (a2048) monitoring of data for month of feb11
- 15mar11
alfa stability (using a2010 date) PolA/polB ratios for feb11
- 15mar11 tiedown
tensions for feb11
- 15mar11 azimuth
motor torques vs velocity for feb11
- 14mar11
corrected 800 Mhz az,za offsets in pointing model
- 11mar11 added
stokes calibration results for 800 Mhz rcvr
- 08mar11 800 Mhz rcvr
bandpass, rms/mean, and sefd vs freq.
- 08mar11 rfi from
alfa motor, dome cameras in 800 Mhz rcvr.
- 02mar11 327 rfi
from:alfa motor controller, dome cameras, distomats.
- 02mar11
updated platform cycles to use change in unbalanced moment as
definition.
- 28feb11
azswings at 327Mhz while transmitting a tone from the 12 meter.
- 25feb11
first look at 12meter rfi from inside the pedestal.
- 22feb11
rfi at 2380Mhz goes off when tower transmitter turned off.
- 22feb11
az
swings done 19jan11 looking at 2380 Mhz rfi.
- 22feb11
completed mosaicing of a2415 430 Mhz maps
- 16feb11
alfa beam 2 has a 5.2 sec oscillation from the dewar bias
monitoring.
- 15feb11 added
dds synth page with bill sisk's schematics.
- 11feb11
counting the number of cycles performed on the telescope 2004.5
thru 2010
- 11feb11
measuring stability in the cband,iflo receiver with the shutter
closed.
- 10feb11 fractional
occurrence of rfi for jan11
- 10feb11 Alfa
Tsys for the month of jan11
- 10feb11
alfalfa
(a2010) monitoring of data for month of jan11
- 10feb11
ages (a2048) monitoring of data for month of jan11
- 10feb11
alfa stability (using a2010 date) PolA/polB ratios for jan11
- 10feb11 tiedown
tensions for jan11
- 10feb11 azimuth
motor torques vs velocity for jan11
- 10feb11
pointing errors after telescope beam reinforcement
- 04feb11
servo
performance test results from 12meter acceptance test.
- 02feb11
pdev voltage sampling on psr0301+19 at 327 and 1410 Mhz
- 02feb11
lbw polB baseline ripple seen by a2335.
- 20jan11
rfi appears at 2380 Mhz
- 13jan11 fractional
occurrence of rfi for dec10
- 13jan11 Alfa
Tsys for the month of dec10
- 13jan11
alfalfa
(a2010) monitoring of data for month of dec10
- 13jan11
ages (a2048) monitoring of data for month of dec10
- 13jan11 tiedown
tensions for dec10
- 13jan11 azimuth
motor torques vs velocity for dec10
- 13jan11
2010 yearly plots of wind velocity from platform.
- 13jan11
2010 yearly plots of platform temperature.
2010 (top)
- 15dec10
327 Tsys vs ra,dec and galactic l,b 10may09 thru 13dec10. Plus
TsysB-TsysB
- 13dec10
az
swing shows that 734-740Mhz rfi is stable. FCC block B low
700Mhz Band.
- 09dec10
lbw
warmup/cooldown cycle on 26/28 oct10 caused jump in Tsys.
- 09dec10
reduced 2010 lbw stokes data.
- 08dec10 fractional
occurence of rfi for nov10
- 08dec10
Alfa
Tsys for the month of nov10
- 08dec10
alfalfa
(a2010) monitoring of data for month of nov10
- 08dec10
ages (a2048) monitoring of data for month of nov10
- 08dec10
alfa stability (using a2010 date) PolA/polB ratios for nov10
- 08dec10
tiedown
tensions for nov10
- 08dec10
azimuth motor torques vs velocity for nov10
- 07dec10
A2335 low interim correlator power levels did not increase Tsys.
- 06dec10
A2335
switching to 750 narrowband IF caused power adjust problems.
- 03dec10 added links
to articles showing FCC 700 Mhz plan and auctions
- 03dec10
reduced 2008/9 cband stokes data.
- 06dec10
switching
to 750 narrowband IF caused power adjust problems.
- 01dec10 added
FCC codes for puerto rico
- 29nov10 added
cal dipole manual (thanks to dana..)
- 29nov10
azimuth dependence of 800 Mhz rcvr rfi.
- 29nov10
hilltop rfi monitoring. added dynamic spectra for last 8 days.
- 27nov10
reduced 2009 lbw stokes data.
- 24nov10
734.7-739.3 rfi in 800Mhz rcvr.
- 23nov10
orion weather station resets accumulated data on 04nov10.
- 22nov10 cband
Tsys increases when horn gortex fabric cover comes loose.
- 22nov10 pointing
errors and 1,2,3 az fits for oct,nov10 after u2-u2l3 beam
repair.
- 20nov10 a2471
data processed... bursts at cband.
- 15nov10
added
46.8 Mhz transmitter control info.
- 12nov10 a2584
cband 1 ghz scan data.
- 10nov10 fractional
occurrence of rfi for oct10
- 10nov10 Alfa
Tsys for the month of oct10
- 10nov10
alfalfa
(a2010) monitoring of data for month of oct10
- 10nov10
ages
(a2048) monitoring of data for month of oct10
- 10nov10
alfa stability (using a2010 date) PolA/polB ratios for oct10
- 10nov10 tiedown
tensions for oct10
- 10nov10 azimuth
motor torques vs velocity for oct10
- 01nov10 started
planetary radar software page.
- 26oct10
12 meter communication timing.
- 20oct10
scope image of protect turnoff timing for the 430 ch xmter.
- 18oct10
tdkips needed to move platform down 1 inch.
- 14oct10
plot a2335 data for 01oct10 thru 13oct10
- 11oct10
monitor lbw stability with 3 1800 sec drift scans.
- 07oct10
a2335 sees instabilities in polA and polB of lbw.
- 09oct10 fractional
occurrence of rfi for sep10
- 09oct10 Alfa
Tsys for the month of sep10
- 09oct10
alfalfa
(a2010) monitoring of data for month of sep10
- 09oct10
ages
(a2048) monitoring of data for month of sep10
- 09oct10
alfa stability (using a2010 date) PolA/polB ratios for sep10
- 09oct10 tiedown
tensions for sep10
- 09oct10 azimuth
motor torques vs velocity for sep10
- 07oct10
430 transmitter transmit pulse timing and constraints.
- 06oct10
alfa
now included in the daily tsys plots. Also generated pdf files.
- 06oct10
430 transmitter: 1 Mhz chirp in tx samples.
- 01oct10
checking 1340 birdie after huffman box gasket changed.
- 30sep10
a2524 800 Mhz rcvr dps processing shows larger ripple in polA.
- 29sep10
4200 second integration lbw to check for stability.
- 28sep10 gps
l3 on for 26 minutes in a2010 data.
- 28sep10
1340 tone coming from radar blanker 1340 Mhz lo.
- 28sep10 added td daily
check output: position,dac values, rpms, torques.
- 27sep10
check stability 430gr receiver. dynamic spectra, avg spc,
rms/mean, pwr vs time.
- 27sep10
scan lbw bands: dynamic spectra, avg spc, rms/mean and total pwr
vs time.
- 27sep10
a2415 430 mapping.
- 26sep10
800 Mhz calibration run after reinstallation.
- 24sep10 5670 cband rfi
8 mhz wide. telescope and hilltop monitor.
- 22sep10 talk living with
rfi at AO for puerto rican spectrum users group added.
- 22sep10 rfi
hilltop monitoring page.. added pictures of equipment on hill.
- 19sep10 fractional
occurrence of rfi for aug10
- 19sep10 Alfa
Tsys for the month of aug10
- 19sep10
alfalfa
(a2010) monitoring of data for month of aug10
- 19sep10
ages
(a2048) monitoring of data for month of aug10
- 19sep10
alfa stability (using a2010 date) PolA/polB ratios for aug10
- 19sep10 tiedown
tensions for aug10
- 19sep10 azimuth
motor torques vs velocity for aug10
- 16sep10
mock narrow band spectra showing birdies close to DC.
- 14sep10
wapps beam5b jumps in power level (attenuators?)
- 10sep10
added platform beam reinforcement page.
- 09sep10
a2335
instabilities in lbw polb seen jul10.
- 05sep10 added
page documenting svn repositories, working directories for code.
- 02sep10
byu phased array sensitivty from single dipole. absorber and
crosses.
- 01sep10
wind at AO during hurrican earl.
- 23aug10
byuPhasedAr Beam A0-A6 position errors 19aug10.
- 19aug10 fractional
occurrence of rfi for jul10
- 19aug10 Alfa
Tsys for the month of jul10
- 19aug10 tiedown
tensions for jul10
- 19aug10
ages
(a2048) monitoring of data for month of jul10
- 19aug10 azimuth
motor torques vs velocity for jul10
- 16aug10
byuPhasedAr Beam A1-A6 position offsets correcting for pointing
errors.
- 13aug10
added monitoring of pdev hardware: fan, temps,voltages
- 11aug10
Mock spectrometer jumps in polA when in hi-res mode.
- 08aug10
added page describing pdev spectrometer diagnostics.
- 04aug10
430 klystron output of 1Mhz chirped signal.
- 08jul10 fractional
occurrence of rfi for jun10
- 08jul10 Alfa
Tsys for the month of jun10
- 08jul10 tiedown
tensions for jun10
- 08jul10
ages
(a2048) monitoring of data for month of jun10
- 08jul10 azimuth
motor torques vs velocity for jun10
- 01jul10
added page to monitor pdev hardware (temps,fans,voltages)
- 30jun10 added page
describing crontab file that makes the daily plots.
- 26jun10
added
byu phased array page.
- 26jun10
updated byu phased array positions in focal plane and on sky.
- 09jun10 fixed
za sin(3az) typo when installing model updates 04jun10
- 08jun10
testing the byu offsets equations using the turret motion.
- 06jun10 fractional
occurrence of rfi for may10
- 06jun10 Alfa
Tsys for the month of may10
- 06jun10
alfalfa
(a2010) monitoring of data for month of may10
- 06jun10
alfa stability (using a2010 date) PolA/polB ratios for may10
- 06jun10 tiedown
tensions for may10
- 06jun10
ages
(a2048) monitoring of data for month of may10
- 06jun10 azimuth
motor torques vs velocity for may10
- 04jun10
added plots of position offsets for byu feed in focal plane and
on sky.
- 04jun10
update pointing models with 1,2,3 az fits to x102 data.
- 28may10
a2463. added plots for 27,28may10. listed bands with
incorrec freqr 18-2may10
- 24may10
spectra jumps at power of 2 channels (23apr10 data)
- 24may10
mock spectrometer. test decimations 1-20.
power,voltage,jumps,bandpass shape.
- 22may10 fractional
occurrence of rfi for apr10
- 22may10 Alfa
Tsys for the month of apr10
- 22may10
alfalfa
(a2010) monitoring of data for month of apr10
- 22may10
alfa stability (using a2010 date) PolA/polB ratios for apr10
- 22may10 tiedown
tensions for apr10
- 22may10
ages
(a2048) monitoring of data for month of apr10
- 22may10 azimuth
motor torques vs velocity for apr10
- 21may10
amman & whitney non-linear analysis of cables ok's cables
for full range of motion.
- 21may10
a2463 w51 on/off-1
- 19may10 a2463
spectral jumps with power of 2 spacing.
- 16may10
5.8 magnitude earthquake 34 km from AO. Td tension variation.
- 11apr10
added documetation for masdmp,wasdmp to dump mock and wapp
aofits files.
- 11apr10
added trackingtime routine to compute lst rise,transit,set as
well as time in the keyhole
- 30apr10
added amman & whitney 29apr10 meeting presentation to limits
page.
- 30apr10
testing the byumap pointing routine on a 3.5 jy src using lbw.
- 27apr10
included plot showing allowable dome za range vs temperature
- 27apr10
added 27apr10 stress table from amman & whitney modeling
- 24apr10
added time on source for byu sources.
- 15apr10
new operational limits for the telescope from ammann and
whitney. (added minKips vs za)
- 11apr10 fractional
occurrence of rfi for mar10
- 11apr10 Alfa
Tsys for the month of mar10
- 11apr10
alfalfa
(a2010) monitoring of data for month of mar10
- 11apr10
alfa stability (using a2010 date) PolA/polB ratios for mar10
- 11apr10 tiedown
tensions for mar10
- 11apr10 azimuth
motor torques vs velocity for mar10
- 04apr10
Cima fixed azimuth drift pattern does not perfom as expected at
ao Declination.
- 31mar10 updated
source tracks by declination to include 1deg steps in za.
- 30mar10
computing the unbalanced moments in the azimuth arm.
- 27mar10
rfi with 1.2 hz periodicity in lbw is from the crosshead.
- 24mar10
mock single pixel observing specifications.
- 19mar10
added
linux datataking programs page.
- 17mar10
The range of za for dome and ch after repairing the broken beam.
- 17mar10
dome za that balances the azimuth arm with ch at stow (8.835)
- 16mar10
document aosock command. Socket i/o from tclInter in vxWorks.
- 16mar10 fractional
occurrence of rfi for feb10
- 16mar10 Alfa
Tsys for the month of feb10
- 16mar10 tiedown
tensions for feb10
- 16mar10 azimuth
motor torques vs velocity for feb10
- 10mar10
tiedown tensions vs az,za, and temp using 08-09 data.
- 09mar10
Repair of broken beam on the platform mar2010
- 07mar10
Dome weight page.
- 28feb10
added
aeronomy
programs page. clp processing.
- 28feb10
clp
computational complexity
- 26feb10
added patriot 12 meter telescope page.
- 23feb10
byu
phased array page.
- 22feb10 fit
tiedown tensions using 2008,2009 data.
- 18feb10 notes for jim
gould.
- 12feb10
updated network switches to show 10Gbit switches.
- 12feb10 fractional
occurrence of rfi for jan10
- 12feb10 Alfa
Tsys for the month of jan10
- 12feb10
alfalfa
(a2010) monitoring of data for month of jan10
- 12feb10
alfa stability (using a2010 date) PolA/polB ratios for jan10
- 12feb10
ages (a2048) monitoring of data for month of jan10
- 12feb10 tiedown
tensions for jan10
- 12feb10 azimuth
motor torques vs velocity for jan10
- 10feb10
tension in tiedown cables during the 12jan10 haitian quake.
- 10feb10
p1693 327 receiver. 3 minute drift scan: dynamic and avg
spectra.
- 08feb10
Converting positionSwitch data to Kelvins. C348,lwide, and
mocks.
- 05feb10
dynamic
loading of tiedowns from telescope motion.
- 03feb10
azimuth motion failures. amplifiers with dc offsets in speed
output.
- 12jan10 calibration
run on BYU 800 MHz antenna
- 11jan10 fractional
occurrence of rfi for dec09
- 11jan10 Alfa
Tsys for the month of dec09
- 11jan10
alfalfa
(a2010) monitoring of data for month of dec09
- 11jan10
alfa stability (using a2010 date) PolA/polB ratios for dec09
- 11jan10
ages (a2048) monitoring of data for month of dec09
- 11jan10 tiedown
tensions for dec09
- 11jan10 azimuth
motor torques vs velocity for dec09
- 11jan10
2009 yearly plots of wind velocity from platform.
- 11jan10
2009 yearly plots of platform temperature.
2009 (top)
- 16dec09
Stokes V sign is flipped if frequency band is inverted.
- 16dec09
disc i/o rates for pdevRaid with 5*1.5TB discs.
- 13dec09
rms baseline degradation of on,off position switching with
position offsets.
- 13dec09 fractional
occurrence of rfi for nov09
- 13dec09 Alfa
Tsys for the month of nov09
- 13dec09
alfalfa
(a2010) monitoring of data for month of nov09
- 13dec09
alfa stability (using a2010 date) PolA/polB ratios for nov09
- 13dec09
ages (a2048) monitoring of data for month of nov09
- 13dec09 tiedown
tensions for nov09
- 13dec09 azimuth
motor torques vs velocity for nov09
- 09dec09
table of radar intermods. added wapp first mixer results.
- 30nov09
Alfa outer beam position errors with mock cimafits and psrfits
files.
- 20nov09
2.5 Mhz oscillations in alfa from bias box.
- 19nov09
alfa total power stability with leds on.
- 19nov09
alfa recover with the leds installed
- 18nov09 fractional
occurrence of rfi for oct09
- 18nov09 tiedown
tensions for oct09
- 18nov09 azimuth
motor torques vs velocity for oct09
- 13nov09
sine wave into echotek card to check timing,frequency.
- 10nov09
converting
spectrometer counts to kelvins with the cals.
- 08nov09 1375 Mhz
rfi at lband.
- 05nov09
calibration scans alfa with alfa rotation angle measurement.
- 01nov09 added
nvidia gpu page
- 26oct09
rdev coded longpulse data from carriage house. plasmalines.
- 21oct09
rdev sine wave test and topside data from 19oct09
- 20oct09
average bandpass for 800 Mhz receiver. data:12oct09
- 12oct09
added psrfits single pixel test dataffiles in stokes mode.
- 07oct09
cima online gui software. how svn is used to maintain it.
- 06oct09 fractional
occurrence of rfi for sep09
- 06oct09 Alfa
Tsys for the month of sep09
- 06oct09
alfalfa
(a2010) monitoring of data for month of jul09
- 06oct09
alfa stability (using a2010 date) PolA/polB ratios for sep09
- 06oct09 tiedown
tensions for sep09
- 06oct09 azimuth
motor torques vs velocity for sep09
- 01oct09
24sep09 rdev test (tamarra's box) with 430 dome and coded long
pulse
- 30sep09
added software packages page: fftw, livedata, predict so far.
- 28sep09
fftw performance on summer 1 cpu (xeon 5420 dual quadcor)
- 17sep09
800Mhz rcvr polB/polA gain,tsys, and sefd ratios
jul09->14sep09
- 14sep09
11sep09 1390 rdr rfi appears in alfalfa data.
- 14sep09 fractional
occurrence of rfi for aug09
- 14sep09 Alfa
Tsys for the month of aug09
- 14sep09
alfalfa
(a2010) monitoring of data for month of jul09
- 14sep09
alfa stability (using a2010 date) PolA/polB ratios for aug09
- 14sep09
ages (a2048) monitoring of data for month of aug09
- 14sep09 tiedown
tensions for aug09
- 14sep09 azimuth
motor torques vs velocity for aug09
- 07sep09 pdev
maintenance procedures: replacing spectrometer box.
- 26aug09
aug09 lro 16,17,20,21 aug. positions and Stokes V of lro xmit.
- 25aug09
az
dependence
tv chn67 after turn off
- 24aug09
ran pdev raid disc i/o test.. pdevs6 raid is slow..
- 23aug09 fractional
occurrence of rfi for jul09
- 23aug09 Alfa
Tsys for the month of jul09
- 23aug09
ages (a2048) monitoring of data for month of jul09
- 23aug09 tiedown
tensions for jul09
- 23aug09 azimuth
motor torques vs velocity for jul09
- 14aug08
800 Mhz cals measured 13aug09 with new Tsys values.
- 13aug09 800
Mhz rcvr. tv chan 67 (789.25 Mhz) turns off.
- 11aug09
sband console rfi in 327 rcvr fixed 11 jun09.
- 10aug09 beam
squint in the 800 Mhz rcvr.
- 10aug09
a2488 800 Mhz dps shows polB rms > polA rms.
- 10aug09
430 dome polA is unstable.
- 09aug09
rfi from azimuth encoders.
- 08aug09
Azimuth encoder1 errors 27jul09->04aug09
- 08aug09
pointing error that occurs during azimuth encoder1 failures
- 30jul09 fractional
occurrence of rfi for jun09
- 30jul09 Alfa
Tsys for the month of jun09
- 30jul09
ages (a2048) monitoring of data for month of jun09
- 30jul09 tiedown
tensions for jun09
- 30jul09 azimuth
motor torques vs velocity for jun09
- 29jul09
added 800 Mhz rcvr to system performance page.
- 03jul09
800 mhz calibration runs with cooled dewar.
- 02jul09
tsys vs freq cband, sbh over entire band.
- 29jun09
CalOn,offs. Averaging over the band vs fitting for wide bands.
- 18jun09 800 mhz
rcvr. freq hop 750/753.7. source of rfi.
- 17jun09
800 Mhz rcvr. ripple appears in polB.
- 17jun09
mock,pdev minimum integration times vs fftlen
- 16jun09
calibration run 800 Mhz rcvr with room temp dewar.
- 14jun09 az swing
800 Mhz rcvr looking at 700-800 Mhz.
- 13jun09
first look at 800 Mhz feed with tv stations off.
- 10jun09 fractional
occurrence of rfi for may09
- 10jun09 Alfa
Tsys for the month of may09
- 10jun09
alfalfa
(a2010) monitoring of data for month of may09
- 10jun09
ages (a2048) monitoring of data for month of may09
- 10jun09 tiedown
tensions for may09
- 10jun09 azimuth
motor torques vs velocity for may09
- 22may09
pdev raid disc i/o speed page
- 16may09
intermods seen in alfa/pdev
- 16may09
table of radar intermods created in alfa/pdev
- 14may09
rfi from lbw crosshead. data taken 05may09
- 10may09 fractional
occurrence of rfi for apr09
- 10may09 Alfa
Tsys for the month of apr09
- 10may09
alfalfa
(a2010) monitoring of data for month of apr09
- 10may09
alfa stability (using a2010 date) PolA/polB ratios for apr09
- 10may09
ages (a2048) monitoring of data for month of apr09
- 10may09 tiedown
tensions for apr09
- 10may09 azimuth
motor torques vs velocity for apr09
- 08may09
lbw dewar warmup time.
- 07may09
Tsys 327 vs sky position jun08 -> may09
- 26apr09
rfi that looks like radar in alfa at 1403, 1424, 1475 Mhz..
- 24apr09
noise bump at 439 Mhz in 430 dome spectra when Tx Hv on.
- 20apr09
visiting committee talk.
- 13apr09
2 az swings with 800 Rcvr centered at 820Mhz.
- 12apr09
az motor 52 fail at az=326 on 11/12apr09
- 11apr09
gps l3 in a2010 data jan09-mar09
- 11apr09
low power tv stations seen in the 700-800 Mhz band.
- 08apr08
2 azswings with 800 rcv @760Mhz. chan62/50 peak values.
- 07apr09 fractional
occurrence of rfi for mar09
- 07apr09 Alfa
Tsys for the month of mar09
- 07apr09
alfalfa
(a2010) monitoring of data for month of mar09
- 07apr09
alfa stability (using a2010 date) PolA/polB ratios for mar09
- 07apr09
ages (a2048) monitoring of data for month of mar09
- 07apr09 tiedown
tensions for mar09
- 07apr09 azimuth
motor torques vs velocity for mar09
- 06apr09
p2445 reports unstable power levels when using the 430 Mhz
receiver
- 06apr09
430 dome spectrum at output of the dewar.
- 03apr09
azswings looking for birdies at 430 Mhz.
- 02apr09
planned usage of the 700 MHz band by emergency services.
- 20mar09
1422.5 Mhz birdie from channel 54.
- 19mar09
GPS L2, GLONASS L2, and compass E6 rfi in 1200-1300 MHz.
- 17mar08
1384,1424 birdies/tones seen in alfa.
- 11mar09
4.7 magnitude earthquake caused tension in td8 to vary.
- 10mar09
collect gps info/measurements to a single page.
- 10mar09
Gain vs frequency. Surface errors from the ruze curves.
- 10mar09 fractional
occurrence of rfi for feb09
- 10mar09 Alfa
Tsys for the month of feb09
- 10mar09
alfalfa
(a2010) monitoring of data for month of feb09
- 10mar09
alfa stability (using a2010 date) PolA/polB ratios for feb09
- 10mar09
ages (a2048) monitoring of data for month of feb09
- 10mar09 tiedown
tensions for feb09
- 10mar09 azimuth
motor torques vs velocity for feb09
- 09mar09
telescope performance and datataking. ausac09 ppt presentation.
- 05mar09
updates to sband transmitter page.
- 01mar09
1385 Mhz birdie in alfa from aerostat. data from 22-23feb09.
- 23feb09
added page with psrfits datafiles that have been taken for
testing.
- 23feb09
pdevs1. run scp while taking data to see %cpu used.
- 22feb09 added
alfa monitoring of dewar temps,drainV,drainI,gateV
- 20feb09 arecibo
radio telescope talk for teachers workshop.
- 18feb09
s2462 sband radar test 30jan09 with satellite orbiting the moon.
- 16feb09
pdev group 0,group 1 power ratios after 1db pads installed.
- 08feb09 fractional
occurrence of rfi for jan09
- 08feb09 Alfa
Tsys for the month of jan09
- 08feb09
alfalfa
(a2010) monitoring of data for month of jan09
- 08feb09
alfa stability (using a2010 date) PolA/polB ratios for jan09
- 08feb09
ages (a2048) monitoring of data for month of jan09
- 08feb09 tiedown
tensions for jan09
- 08feb09 azimuth
motor torques vs velocity for jan09
- 04feb09
spectral scans with 800 Mhz rfi feed taken 18dec08
- 27jan09
psrfits2fil: convert pdev psrfits format to sigproc filterbank
format.
- 20jan09
locations for 1db pads in pdev to equalize grp0,grp1 power
levels.
- 20jan09
added some documentation on pdev adjusting the power levels.
- 18jan09
gpsL3 test baseband data taken 23sep08.
- 16jan09 fractional
occurrence of rfi for dec08
- 16jan09 Alfa
Tsys for the month of dec08
- 16jan09
alfalfa
(a2010) monitoring of data for month of dec08
- 16jan09
alfa stability (using a2010 date) PolA/polB ratios for dec08
- 16jan09
ages (a2048) monitoring of data for month of dec08
- 16jan09 tiedown
tensions for dec08
- 16jan09 azimuth
motor torques vs velocity for dec08
- 14jan09
updated info on alfa motor controller page.
2008 (top)
- 16dec08
instructions for daily compressing of a2130 (galfacts) data.
- 15dec08
lbw cals remeasured 08dec08. installed 15dec08
- 14dec08 fractional
occurrence of rfi for nov08
- 14dec08 Alfa
Tsys for the month of nov08
- 14dec08
alfalfa
(a2010) monitoring of data for month of nov08
- 14dec08
alfa stability (using a2010 date) PolA/polB ratios for nov08
- 14dec08
ages (a2048) monitoring of data for month of nov08
- 14dec08 tiedown
tensions for nov08
- 14dec08 azimuth
motor torques vs velocity for nov08
- 12dec08 update
utc-ut1 for leap second on 01jan09
- 10dec08
gregorian motor torques during galfacts scanning
- 07dec08
monitoring pdevsN server cpu/disc with iostat.
- 05dec08
pdev spectrometer: setting the registers for 16 bit sampling.
- 02dec08
pdev running galfacts can't keep up data rate at 65
Mb/sec.
- 27nov08
trying to remove the 1400Mhz lbw resonance in software
- 19nov08
a2130 .. added list of files recorded..
- 18nov08
pdev winking cal occasionally starts with a phase offset.
- 18nov08
a2130 galfacts survey project page. info about
files,compression..
- 09nov08 fractional
occurrence of rfi for oct08
- 09nov08 Alfa
Tsys for the month of oct08
- 09nov08
alfalfa
(a2010) monitoring of data for month of oct08
- 09nov08
alfa stability (using a2010 date) PolA/polB ratios for oct08
- 09nov08
ages (a2048) monitoring of data for month of oct08
- 09nov08 tiedown
tensions for oct08
- 09nov08 azimuth
motor torques vs velocity for oct08
- 30oct08
Check how clean 839 Mhz is using hilltop monitoring data for
oct08.
- 28oct08 wide
spectral variations in upper half of ch band.
- 06oct08 fractional
occurrence of rfi for sep08
- 06oct08 Alfa
Tsys for the month of sep08
- 06oct08
alfalfa
(a2010) monitoring of data for month of sep08
- 06oct08
alfa stability (using a2010 date) PolA/polB ratios for sep08
- 06oct08
ages (a2048) monitoring of data for month of sep08
- 06oct08 tiedown
tensions for sep08
- 06oct08 azimuth
motor torques vs velocity for sep08
- 01oct08
psrfits test program to time cfitsio i/o rates.
- 29sep08
sbw cal values updaed from 26sep08 cal measurements.
- 26sep08
lbw cals updated from 26sep08 cal measurements.
- 26sep08
alfa long recovery times vs alfa rotation angle. data taken
24sep08.
- 24sep08
pdev adc blanking of faa, remy 1290 radar. az=180, za=17 to 17.5
- 23sep08
alfa long recovery times after bias changes (on sky)
- 21sep08
check alfa stability after bias voltages changed.
- 18sep08 added idl
routines for satellite predection at AO.
- 18sep08
alfa long recovery. modify 3a bias board. Change V,I bias for 3A
and 3B
- 16sep08
alfa long recovery. Remove zener 3b, amp3 bias circuit. Modify
Vd,I
- 16sep08
picture showing alfa, pdev layout.
- 12sep08
Modifying the bias currents with pulsed noise into the alfa
horn.
- 05sep08
az strips with the 800 Mhz rfi feed looking for turret position.
- 05sep08 fractional
occurrence of rfi for aug08
- 05sep08 Alfa
Tsys for the month of aug08
- 05sep08
alfalfa
(a2010) monitoring of data for month of aug08
- 05sep08
alfa stability (using a2010 date) PolA/polB ratios for aug08
- 05sep08
ages (a2048) monitoring of data for month of aug08
- 05sep08 tiedown
tensions for aug08
- 05sep08 azimuth
motor torques vs velocity for aug08
- 04sep08
22-28aug08 gps sv26 flies thru beam during proj a2010.
- 02sep08
measure pdev pfb overflow vs a/d rms and fftlen using a sine
wave input.
- 29aug08 lbw
cal values meaured and installed.
- 28aug08 fractional
occurrence of rfi for jul08
- 28aug08 Alfa
Tsys for the month of jul08
- 28aug08
ages (a2048) monitoring of data for month of jul08
- 28aug08 tiedown
tensions for jul08
- 28aug08 azimuth
motor torques vs velocity for jul08
- 27aug08
Pulsed noise into alfa horn shows long recovery is in dewar.
- 16jul08
tsys 327 vs date,hr,ra,dec
- 10jul08
pdev spectral bandpass shape and spectral jumps with low
a/drms
- 04jul08 fractional
occurrence of rfi for jun08
- 04jul08
Alfa
Tsys for the month of jun08
- 04jul08
ages (a2048) monitoring of data for month of jun08
- 04jul08
tiedown
tensions for jun08
- 04jul08
azimuth motor torques vs velocity for jun08
- 02jul08
lbw cooldowns.. omt stage not getting below 180K
- 25jun08
faa radar cause pdev a/d to overflow causing spectral artifacts.
- 25jun08
intermods/harmonics created when clipping a sine wave.
- 24jun08
modeling the faa radar total power period for pulsar observers.
- 24jun08 pdev
spectrometer added list of datafiles taken by month.
- 23jun08
added tiedown loadcell locations and serial numbers.
- 23jun08
SEFDA/SEFDB for 327 rcvr shows TsysB/TSYSA=1.1
- 23jun08
added 327 dewar to receiver monitoring.
- 23jun08
11jun08 mot82 azTq reads nonzero when power off.
- 22jun08
fit 1az, 3az term to pointing runs to update model.
- 19jun08
pdev take 12 secs of data with galfacts setup (sitting 270,10).
- 15jun08
pdev time domain sampling does not have an 8bytes record header.
- 14jan08
pdev spectrometer specifications.
- 13jun08
added cal values for 327 cooled receiver.
- 13jun08
pdev spectrometer. spectral bandpass, rms vs Ashift param
for 16 bit sampling.
- 12jun08
more a1300 calibrate oh spectra for tom..
- 10jun08 fractional
occurrence of rfi for may08
- 10jun08
Alfa
Tsys for the month of may08
- 10jun08
alfalfa
(a2010) monitoring of data for month of may08
- 10jun08
ages (a2048) monitoring of data for month of may08
- 10jun08
alfa stability (using a2010 date) PolA/polB ratios for may08
- 10jun08
tiedown
tensions for may08
- 10jun08
azimuth motor torques vs velocity for may08
- 09jun08
td 12 encoder interface board fails from lightning strike.
- 09jun08
added system performance by month page.
- 08jun08
started
page with pdev fpga code
- 06jun08
added pdev link from home page
- 30may08
a1300 calibrate oh spectra for tom..
- 29may08
tiedown positions during the painting project.
- 24may08
platform, tiedown motion during vlbi cal scan 15may08
- 21may08
x102
run
output
from
20may08
cband.
- 19may08
a2327 dataprocessing.
- 16may08
process a2228 xband data looking for O2.
check xband stability.
- 05may08 fractional
occurrence of rfi for apr08
- 05may08
Alfa Tsys for the month of apr08
- 05may08
alfalfa (a2010) monitoring of data for month of apr08
- 05may08
ages (a2048) monitoring of data for month of apr08
- 05may08 alfa
stability (using a2010 date) PolA/polB ratios for apr08
- 05may08 tiedown
tensions for apr08
- 05may08
azimuth motor torques vs velocity for apr08
- 03may08
pdev spectral bandpass and rms vs ashift parameter for 8 and 16
bit packing.
- 02may08
gain and bandpass shape of the digital low pass filters of pdev.
- 02may08
a2390.. hi cal not firing at end of scan.
- 01may08
when faa radar broadcasts at 1330,1350.. 2008 thru april
- 26apr08
bbmixer compression levels from a sine wave input.
- 24apr08
pdev
A/D
rms
values
vs
mixerGain
and
compression
levels.
On
sky
and
noisesrc.
- 22apr08
lbw polb is noisier than polA.
- 21apr08
pnetctl information
- 08apr08
07apr08 td 8 stops moving. commandvel=0 and level jmps in torque
readback.
- 06apr08 fractional
occurrence of rfi for mar08
- 06apr08
Alfa Tsys for the month of mar08
- 06apr08
alfalfa (a2010) monitoring of data for month of mar08
- 06apr08
ages (a2048) monitoring of data for month of mar08
- 06apr08 alfa
stability (using a2010 date) PolA/polB ratios for mar08
- 06apr08 tiedown
tensions for mar08
- 06apr08
azimuth motor torques vs velocity for mar08
- 05apr07
pulsed power injected into alfa down converted to see long
revcovery times.
- 05apr08
alfa if/lo system
- 25mar08
The za locations were alfa power jumps occur when at the
meridian.
- 25mar08
alfa power jumps with long recovery times using pdev
- 17mar08
pdev spectral bandpass shape vs A/D input level and PSHIFTvalue.
- 14mar08
pdev bbm signal levels. Injecting a sine wave at the bbm input.
- 14mar08
input signal levels to generate a full scale a/d output for
pdev.
- 10mar08
alfa power jump and long recovery time from radars using pdev.
- 09mar08 fractional
occurrence of rfi for jan08,feb08
- 09mar08
Alfa Tsys for the months jan08,feb08
- 09mar08
alfalfa (a2010) monitoring of data for month of
jan08,feb08
- 09mar08
ages (a2048) monitoring of data for month of
jan08,feb089
- 09mar08
alfa
stability (using a2010 date) PolA/polB ratios for jan08,feb08
- 09mar08 updated
system performance page for jan08,feb08
- 08mar08
rdev filter shapes with 17 and 11 db roll off coefficients.
- 21feb08
az motor torques for motor group 1 and motor group 2.
- 20feb08
Added iflo block diagrams.
- 07feb08
check lbw sefdpolA/sefdPolB after adjusting the dewar biases.
- 06feb08
first start at params needed to config pdev from cima.
- 06feb08
az motor torques after amplifier pots adjusted.
- 05feb08
check lbw polB Tsys using sefd taken 18dec07
- 05feb08
spread in az motor torques jan08.
- 04feb08
added monthly monitoring of azimuth motor torques.
- 04feb08
added jon hagens vertex memos.
- 04feb08
check td12 after adjusting the current limit on amp.
- 21jan08
added table of filter programs to make ascii dumps of data and
monitor files.
- 18jan08
pointing model gradients
- 18jan08
td 12 has been stopping off and on. Torque limit set too low.
- 11jan08 fractional
occurrence of rfi for dec07
- 11jan08
Alfa Tsys for the month of dec07
- 11jan08
updated system performance page for dec07
- 11jan08
updated wind monitor page with yr 2007 summary
- 11jan08
updated platform temp page with yr 2007 summary
- 11jan08
added temperature sensor page in hardware devices.
- 09jan08
vertex motor, amp, compboard settings.
2007 (top)
- 16dec07
pointing model 17 after the painting project.
- 15dec07
status recommissioning telescope after the painting..
- 11dec07
plotted az motor torques from az swings. Motor 81's torque is
too high.
- 11dec07
az bending errors.
- 10dec07
Moved td ref position using 10dec07 data then redid az swing to
check 1az term.
- 10dec07
az
swings measuring platform tilt.
- 09dec07
equalizing the tiedown positions after removing the lead in the
platform corners.
- 07dec07
remeasure platform tilt using tilt sensors after adjusting with
06dec07 data.
- 06dec07
measure platform tilt with tilt sensors after painting with
current td positions.
- 04dec07
more updates on aofits header.
- 27nov07
refit dut1 (utc-ut1) using data 27nov07 to 22nov08.
- 26nov07
notes on aofits header.
- 16nov07
added ntp entry to hardware page to keep track of ntpd
measurements.
- 30oct07
pdev mixer i,q phase relation. added loband data with lo at 205
and 215
- 19oct07
added platform motions page to group all the platform motion
measurements.
- 12oct07
setting the gain levels in the pdev spectrometer.
- 11oct07
simulating the fir portion of the polyphase filter bank for
pdev.
- 10oct07
checking spectral output vs pdev dshift param.
- 01oct07
converting spectrum back to rms volts (notes to myself)
- 28sep07
simulating the gain of the digital low pass filters.
- 26sep07
bump in center of bandpass from mixture chassis stable in time.
- 26sep07
location of noise rms bits in DLPF processing (setting
hr_shift).
- 20sep07
spectral increase +/- 3 MHz about center is gain related.
- 19sep07
noise spectra with and without the mixer chassis.
- 18sep07
pdev mixer bandpass using noise source.
- 29aug07
testing time domain dumping on the pdev spectrometer.
- 23aug07
monitor programs: montel, monpnt, montie, etc.. now run on linux
- 06aug07
Added how jmocks routines start,run, stop in pdev software
documentation.
- 06aug07
recomputed lbw stokes plots using carl's latest and greatest
fix.
- 04aug07
added software page for pdev software. What i'm doing and how to
rebuild jeffs system.
- 04jul07
Pdev device. Looking at birdies close to dc with radar fpga
(with 25Mhz filter).
- 01jul07
first look at pdev mixer chassis bandpass shape.
- 29jun07
computed
stokes calibration data lbw for jan2006->jun07
- 21jun07
pdev accpetance testing. Updated with interim correlator
comparison. Include on/off spectra.
- 14jun07
pdev acceptance testing. included the rest of the boxes. added
rms vs integration time.
- 12jun07
pdev acceptance testing .. include longer integrations.
- 11jun07
p1978 wapp pulsar polarization runs but 2 rather than 1 diodes
used.
- 06jun07
acceptance testing the spectrometer boxes.
- 09may07
Strong 60 hz seen in the 327 receiver 15apr07 to 03may07.
- 07may07
fractional occurrence of rfi for apr07
- 07may07
Alfa Tsys for the month of mar07
- 07may07
alfalfa (a2010) monitoring of data for month of apr07
- 07may07
alfa stability (using a2010 date) PolA/polB ratios for apr07
- 07may07
ages (a2048) monitoring of data for month of apr07.
- 07may07
updated system performance page for apr07
- 03may07
modeling the azimuth pointing errors.
- 02may07
recomputed the azimuth encoder runout (with a simpler method).
- 28apr07
frequency spectrum (cycles in 360 degrees) for the az encoder
runout data.
- 26apr07
24apr07 pointing checks with sbn.
- 25apr07
added pdev spectrometer idl routines for testing
- 25apr07
added rdev radar processor routines for testing.
- 24apr07
23apr07 pointing checks with sbn.
- 22apr07
polA resonance lbw at 1400 Mhz gone after omt bolts tightened.
PolB at 1395 still there.
- 22apr07
21apr07 pointing checks with sbn.
- 21apr07
20apr07 pointing checks with sbn.
- 21apr07
19apr07 pointing checks with sbn.
- 19apr07
included
wind
monitoring
for
the
previous
week.
- 19apr07
18apr07 pointing checks with sbn.
- 18apr07
What happens to the dome when we move the ch from 2 to 20 degs
za.
- 18apr07
17apr07 pointing checks with sbn after alfa lowered.
- 18apr07
13apr07 pointing checks with sbn.
- 18apr07
plot
focus
curves
feb03
and
2006
.
- 15apr07
The rotation angle on the sky when we rotate the platform.
- 13apr07
07apr07
tilt
sensor
azswings,
adjust
td's,
after
lead
removed
from
corner
12.
- 13apr07
1330/1350 faa radar ipps back to normal sequence.
- 13apr07
07apr07 tilt sensor az swings and then readjustment of the
tiedown reference.
- 13apr07
04apr07 tilt sensor azswings, adjust td's, more azswings after
lead removed.
- 12apr07
12apr07 pointing checks with sbn.
- 12apr07
09apr07 pointing checks with lbw after installation in dome.
- 12apr07
11apr07 pointing checks.
- 12apr07
1330/1350 faa radar ipps jump by 9 useconds.
- 11apr07
10apr07 pointing checks.
- 11apr07
azimuth encoder 1 failed twice in the morning.
- 09apr07
encoder 1 fault 08apr07. eventually reset by power dip.
- 08apr07
07apr07 pointing checks after tiedowns reset to correct for
weight removed from corner 12.
- 08apr07
05_06apr07 pointing checks after azimuth encoders reset.
- 07apr07
04apr07 dome local disable cable breaks. Both azimuth encoders
reset.
- 06apr07
04apr07 pointing check after tiedown positions moved.
- 03apr07
radar seen 28mar07 in the serendip 5 spectrometer using alfa.
- 03apr07
02apr07 pointing checks after half the lead from corner 4
lowered.
- 02apr07
01apr07 pointing checks.
- 02apr07
fractional occurrence of rfi for mar07
- 02apr07
Alfa Tsys for the month of mar07
- 02apr07
alfalfa (a2010) monitoring of data for month of mar07
- 02apr07
alfa stability (using a2010 date) PolA/polB ratios for mar07
- 02apr07
updated system performance page for mar07
- 31mar07
31mar07 pointing checks after laser shack removed.
- 31mar07
tilt sensor az swings after laser shack removed.
- 31mar07
30mar07 pointing checks.
- 30mar07
tilt sensor az swings showing 1,3az terms prior to painting
project.
- 29mar07
pointing checks before painting project started
- 08mar07
bad cable wapp 4 brd1 polB (alfa 6B) causes jumps in the gain
- 08mar07
fractional occurrence of rfi for feb07
- 08mar07
Alfa Tsys for the month of feb07
- 08mar07
alfalfa (a2010) monitoring of data for month of feb07
- 08mar07
alfa stability (using a2010 date) PolA/polB ratios for feb07
- 08mar07
updated system performance page for feb07
- 07mar07
added monthly monitoring of AGES a2048 alfa data.
- 06mar07
a2125 continuum mapping at 430 Mhz.
- 05mar07
when the faa radar transmits at 1330 and 1350 Mhz
- 05mar07
Added the faa radar page to the rfi pages.
- 19feb07
lbw beam maps computer from the calibration scans.
- 19feb07
amplifier 21,22 of dome going not ready.
- 17feb07
telescope performance talk for feb07 visiting committe.
- 13feb07
lbw sefd around 1100 Mhz with some rfi plots.
- 09feb07
refit alfa tsys using cal data 01jan05 thru 31jan07.
- 07feb07
gps l3 test on 06dec06 in the a2010 data.
- 06feb07
fractional occurrence of rfi for jan07
- 06feb07
Alfa Tsys for the month of jan07
- 06feb07
alfalfa (a2010) monitoring of data for month of jan07
- 06feb07
alfa stability (using a2010 date) PolA/polB ratios for jan07
- 06feb07
updated system performance page for jan07
- 02feb07
First look at the faa radar with the pdev spectrometer.
- 02feb07
measuring where the pdev specrometer satures (pfb sat.)
- 31jan07
if2 noise source test of pdev spectrometer. Include rms vs
integration time.
- 23jan07
2 minute on/offs with pdev (jeffmocks new spectrometer)
- 20jan07
system performance page: llimit to 2000 pnts,.pdf files added,
old data moved to separate page.
- 18jan07
added ao network switch contents.
- 12jan07
Platform temperature yearly summary plots for 2006.
- 12jan07
Windmeter yearly summary plots for 2006.
- 12jan07
fractional occurrence of rfi for dec06
- 12jan07
Alfa Tsys for the month of dec06
- 12jan07
alfalfa (a2010) monitoring of data for month of dec06
- 12jan07
alfa stability (using a2010 date) PolA/polB ratios for dec06
- 12jan07
updated system performance page for dec06
2006 (top)
- 14dec06
large birdies at 418.625 and 419.2 Mhz.
- 11dec06
update the pointing models for sin(za) and 1 az terms.
- 10dec06
fit new value for UT1-UTC (dec06 thru jun07).
- 06dec06
4 az swings with tilt sensors at za=10 to measure platform
residual tilt.
- 06dec06
Updated angle of tilt sensor on the rotary floor for data taken:
2005,2006.
- 05dec06
fractional occurrence of rfi for nov06
- 05dec06
Alfa Tsys for the month of nov06
- 05nov06
alfalfa (a2010) monitoring of data for month of nov06
- 05dec06
alfa stability (using a2010 date) PolA/polB ratios for nov06
- 05dec06
updated system performance page for nov06
- 28nov06
check the linearity of the programmable gain in the downstairs
IFLO.
- 28nov06
pointing checks nov06 after az encoder error.
- 22nov06
a2125 430Mhz continuum maps of the individual fields.
- 21nov06
How well was encoder 1 realigned after the 14nov06 encoder
failure.
- 12nov06
beam1a of alfa is noiser than the other beams.
- 07nov06
fractional occurrence of rfi for oct06
- 07nov06
Alfa Tsys for the month of oct06
- 07nov06
alfalfa (a2010) monitoring of data for month of oct06
- 07nov06
alfa stability (using a2010 date) PolA/polB ratios for oct06
- 07nov06
updated system performance page for oct06
- 31oct06
cable test lbw in upstairs iflo shows loose cables in 750 mixer
chassis.
- 28oct06
az Encoder board failure 14oct06.
- 18oct06
looking at the rfi from the axis web camera.
- 12oct06
Dewar monitor failure was the upstairs ethernet repeater.
- 12oct06
added dewar monitoring hardware page.
- 09oct06
compare 614 khz comb strength with new and old encoders.
- 09oct06
added az,za encoder page. Include comb at 430 Mhz.
- 05oct06
az,za, time dependence of the 614 khz comb in 327.
- 04oct06
614 Khz comb in 327 Mhz reciever.
- 03oct06
added 327 birdies with azimuth dependence to the rfi frequency
list.
- 03oct06
fractional occurrence of rfi for sep06
- 03oct06
Alfa Tsys for the month of sep06
- 03oct06
alfalfa (a2010) monitoring of data for month of sep06
- 03oct06
alfa stability (using a2010 date) PolA/polB ratios for sep06
- 03oct06
updated system performance page for sep06
- 29sep06
added alfa stablity, jan06 thru jun06.
- 28sep06
added alfa stablity page. Used a2010 data and ratio Pola/PolB
added months aug06,sep06
- 25sep06
12 second radar at 1378 Mhz appears for 430 seconds on 05sep06
- 24sep06
fractional occurrence of rfi for aug06
- 24sep06
Alfa Tsys for the month of aug06
- 24sep06
alfalfa (a2010) monitoring of data for month of aug06
- 24sep06
updated system performance page for aug06
- 31aug06
new lbw cal values installed. measured 06jul06.
- 30aug06
cable sag measurements .. added avg td tensions.
- 29aug06
new cband cal values installed. Measured 06jul06.
- 24aug06
test the rf monitor at the visitors center with the 430
transmitters.
- 20aug06
Added daily monitor of tiedown tensions.
- 17aug06
430 Dome. Fit antenna gain vs frequency.
- 12aug06
check 430yagi performance after the feed was moved closer to the
linefeed.
- 09aug06
noise/gain stability 4-6 ghz mmic, cbh using sqLaw det. and
wapps
- 06aug06
fractional occurrence of rfi for jul06
- 06aug06
Alfa Tsys for the month of jul06
- 06aug06
alfalfa (a2010) monitoring of data for month of jul06
- 03aug06
test stability of mmic amp in cryo lab. also check cbh and noise
src.
- 30jul06
test stability of mmic amp in cryo lab.
- 27jul06
added punta salinas radar page.
- 22jul06
punta salinas radar blanking test.
- 18jul06
az,za dependence of rfi using transmitted birdies.
- 17jul06
1200 Mhz narrow birdie.
- 14jul06
1188 Mhz birdie, 66 Mhz comb from kronos timecard reader.
- 13jul06
cat walk weight survey by felipe soberal
- 13jul06
added lbw sefd fits Mar03 thru jun06.
- 13jul06
added sefd page for receivers
- 13jul06
added sefd description page
- 10jul06
birdie near 1320 Mhz.
- 07jul06
az,za position and tiedown tensions during main cable sag
measurements (mar06)
- 05jul06
fractional occurrence of rfi for jun06
- 05jul06
Alfa Tsys for the month of jun06
- 05jul06
alfalfa (a2010) monitoring of data for month of jun06
- 29jun06
updated system performance page for may06
- 29jun06
alfalfa (a2010) monitoring of data for month of may06
- 29jun06
Alfa Tsys for the month of may06
- 29jun06
fractional occurrence of rfi for may06
- 02jun06
430 yagi jumps are the rfi birdies with 1 minute periodicity
- 01jun06
System performance 430 yagi after capacitor repaired.
- 24may06
moved 430 yagi info from 430ch page to its own page.
- 12may06
Compare the focus curves for the old and new sb tx horn.
- 12may06
sband tx focus curve after the new horn was installed.
- 11may06
fractional occurrence of rfi for apr06
- 11may06
distomat birdies after the windows changed (corrected distomat
numbering).
- 10may06
a1731 looking at the processe datamap,beammap, and uvdata.
- 07may06
performance of new sband tx horn.
- 03may06
distomat birdies after the windows were changed (data taken
07apr06)
- 03may06
updated system performance page for apr06
- 03may06
alfalfa (a2010) monitoring of data for month of apr06
- 03may06
Alfa Tsys for the month of apr06
- 28apr06
sband tx focus curve before the new horn was installed.
- 27apr06
focus curve using sbn rcv on J0745+101 (taken 25apr06).
- 26apr06
trendata ethernet switch rfi with and without fiber active.
- 21apr06
one of alfa's refrigerators fails on 18apr06.
- 21apr06
td12 startup trouble was digital i/o signal conditioning board
- 19apr06
the 750,40 Khz combs in 327 receiver coming from sband console
in dome.
- 16apr06
750 Khz and 39.45 Khz combs in 327 receiver.
- 15apr06
testing the trenddata gigabit switch for rfi.
- 13apr06
added a2127 Xband data for 08mar06 and 03apr06.
- 11apr06
sbn cals meausured at 2380 and 2280..
- 11apr06
stokes calibration sbn new cals using 3C138. data daken 17feb06
- 09apr06
430 ch yagi performance after latest move.
- 08apr06
mars express quick look.
- 06apr06
updated my rfi page. Moved things from my home page to here.
- 05apr06
ntp time jumps on the wapps 2003-2006. What was causing them.
- 03apr06
added things to check when configuring a linux kernel for
datataking.
- 03apr06
When the ntpd on aeron lost sync 2003-2006.
- 03apr06
updated system performance page for mar06
- 03apr06
alfalfa (a2010) monitoring of data for month of mar06
- 02apr06
fractional occurrence of rfi for mar06
- 01apr06
Alfa Tsys for the month of mar06
- 01apr06
echotek configuration causes i,q samples to not have a 90deg
phase difference.
- 29mar06
Document vxWorks programs by vme crate.
- 28mar06
look at new Ac units. compare with older units to see if comb is
gone.
- 27mar06
added radar decoding explanation to my notes.
- 19mar06
linefeed, yagi overlap at 100 km.
- 16mar06
A description of tiedown a/d signals
- 16mar06
td 12 does not always start up when it is commaned to.
- 14mar06
Start mechincal oscillations page.
- 14mar06
Azimuth encoder difference and the az encoder rack gear.
- 13mar06
total power oscillations while tracking the crab at cbh.
- 09mar06
Measure tsys for the 9 filters of lbw.
- 08mar06
Az encoder 2 (ch side) jumps.
- 07mar06
lbw cal polB . tapping on cables external to dewar to not change
cal.
- 07mar06
lbw 1230-1470 filter look to see if Tsys changes as we switch it
in /out.
- 07mar06
The jun05 warmup/cooldown of lbw actually caused resonances in
polA and polB.
- 04mar06
lbw reran stokes analysis for dec05,jan06,feb06 to use new cal
values.
- 04mar06
Alfa Tsys for the month of feb06
- 04mar06
fractional occurrence of rfi for feb06
- 04mar06
alfalfa (a2010) monitoring of data for month of feb06
- 03mar06
td 12 releases for 10 minutes.
- 03mar06
added idl routines to look at tiedown data.
- 02mar06
metmon: idl routine to monitor meteors. added positioning
cabability
- 01mar06
added accessing the cal values page:
- 28feb06
remeasured lbw cal values after 04dec06 polB jump.
- 28feb06
lbw polB cal value jumps by 20% on 04dec05.
- 24feb06
Harmonics created in the alfa/wapp receiver with cfr=1385.
- 24feb06
added idl routine alfawappharm to compute location of harmonics
in alfa/wapp.
- 17feb06
Added Ao serverlist.
- 15feb06 tried to make the different receiver
pages a little more similar.
- 15feb06
put 430 dome cals in cals page (these values were measured back
in 2001).
- 08feb06
Added yearly summary plots of windmeter data.
- 07feb06
The radar blanker multiplexor (calixto's box) needs to be worked
on.
- 05feb06
Daily, yearly platform temperatures now on web.
- 04feb06
Daily windmeter data now on web.
- 03feb06
fractional occurrence of rfi for jan06
- 03feb06
alfalfa (a2010) monitoring of data for month of jan06
- 03feb06
Alfa Tsys for the month of jan06
- 03feb06
updated system performance page for jan06
- 01feb06
Added monitoring of wind meter data.
- 28jan06
check the azimuth encoder offset after the nov05 failure.
- 24jan06
47 Mhz On, off 3C348.0 (Hercules A) Tsys,gain,Beamwidth,sefd
- 24jan05
A first look at the 47 Mhz bandpass, dynamic spectra, rfi.
- 21jan06
rms voltage for optimum sampling with 2,3,4, and 5 bits.
- 18jan06
check the calibration source fluxes.
- 17jan06
compression in the alfa receiver with the 100 Mhz filter
installed.
- 16jan06
28oct05 az swings with 327 showed birdies with 1 minute
periodicities.
- 15jan06
wide band look at ac dome birdies in 327 Mhz reciever.
- 14jan06
3rd harmonics of fm stations in 327 receiver.
- 13jan06
placing a hipass and bandpass filter before the 327Rcvr first
amp.
- 13jan06
check xband receiver after testing in lab.
- 11jan06
re arranged the rfi measurements page (might have broken some
links??)
- 10jan06
1 Mhz comb of Ac units in dome also seen in 430 Mhz rcvr.
- 09jan06
26dec05 lbw no power polB for 1 file (probably downstairs iflo
problem).
- 09jan06
added links to vanvleck correction papers: fred schwab,
kulkarni/heiles.
- 05jan06
Alfa Tsys for the month of dec05
- 05jan06
fractional occurrence of rfi for dec05.
- 05jan06
alfalfa (a2010) monitoring of data for month of dec05
- 05jan06
updated system performance page for dec05
2005 (top)
- 14dec05
Instructions for crossing midnight at the end of the year.
- 13dec05
Started adding list of combs to the rfi freq list page.
- 12dec05
9 level correction problem with wapp pulsar folding.
- 07dec05
fractional occurrence of rfi for nov05.
- 07dec05
alfalfa (a2010) monitoring of data for month of nov05
- 07dec05
Alfa Tsys for the month of nov05
- 06nov05
updated system performance page for nov05
- 05dec05
430 Mhz birdies with 1 minute periodicity.
- 01dec05
bits used in the hp3488 digital output device for the downstairs
IF2.
- 26nov05
trouble with lbw polB filter 3 (1230-1470 Mhz).
- 22oct05
gain curves of all recievers. updated the cband gain curve,
added 327.
- 22oct05
cband gain curves fit to data may05-oct05.
- 22oct05
cband system performance used for gain curve fits
- 21nov05
performance of mathews 430 test feed with cals installed.
- 05nov05
The azimuth dependence of birdies in the 327 Mhz reciever.
- 04nov05
alfalfa (a2010) monitoring of data for month of oct05
- 02nov05
updated system performance page for oct05
- 02nov05
Alfa Tsys for month of oct05.
- 02nov05
fractional occurrence of rfi for oct05.
- 29oct05
lbw calibration jumps for murray's oh data.
- 28oct05
standing waves measured by moving the platform with the
tiedowns.
- 21oct05
The output of the 327 Mhz rcvr after the 1st amp (and before the
filter).
- 17oct05
location of faa 1350 radar aliases in alfa.
- 16oct05
1 Mhz comb in 327 coming from digital control units in dome AC
units.
- 14oct05
Cal stability for mathews 430 feed checked.
- 13oct05
added idl routine atmclp to process coded long pulse data from
ri rawdat.
- 10oct05
some linux kernels see no speed up when running 2 processes on a
dual processor cpu.
- 07oct05
Tsys jump of sbw aug05 is from the cal value changing.
- 06oct05
fractional occurrence of rfi for sep05.
- 06oct05
Measure the 1300/1400 Mhz birdies in lbw. Most comes from the
wapps.
- 05oct05
Alfa Tsys for month of sep05.
- 05oct05
a2010 (alfalfa) monitor data for month of sep05.
- 29sep05
added 430 carriage house cal values.
- 27sep05
atm digital receiver glitches every 100 ipps.
- 26sep05
added idl shsxxx routines to read atm echotek digital receiver
data.
- 20sep05
cell phone harmonics in lbw while transmitting from inside the
dome.
- 20sep05
added cell phone info page.
- 16sep05
td8 analog input problems.
- 15sep05
alfa: the cal deflection is not a function of the feed rotation
angle.
- 14sep05
TsysA-TsysB "polarized Tsys" with the alfa rcvr. Az spins and
feed rotation.
- 13sep05
tilt sensor measurement while rotating turret floor (after
rolling surface section replaced).
- 07sep05
tension in tiedowns after the loadcells were recalibrated.
- 05sep05
fractional occurrence of rfi for aug05.
- 05sep05
The 16Khz comb at 327. Grounding the motor shields helps for
awhile.
- 03sep05
Alfa Tsys for month of aug05.
- 03sep05
Updated system performance or aug05.
- 27aug05
the 16Khz comb at 327 Mhz is coming from the alfa motor
controller.
- 25aug05
Testing the cal and measuring Tsys for mathews 430 dipole on ch.
- 17aug05
Fitting Tsys for alfa using data jan05 thru jul05.
- 14aug05
sband bistatic run pointing problems (17apr05 run)
- 11aug05
installed new cband cal values (measured 10jun05).
- 11aug05
updated system performance for jul05.
- 11aug05
added fractional occurrence of rfi for july05
- 10aug05
Plot Tsys vs za by month for alfa. Added months
mar,apr,may,jun,jul05.
- 08aug05
check performance of cband hi for thankins.
- 08aug05
17apr05: tension in tiedowns during sband run.
- 06aug05
p2030 alfa pulsar search sees large 12 second compression in the
01aug05 data (from faa radar).
- 19jul05
galfa pixels 4a,6a switched 11jul05 thru 19jul05
- 13jul05
power levels drift/jump in pixel2b of galfa spectrometer.
- 11jul05
alfa: replacing pixel 4a fiber receiver makes 174Khz ripple go
away.
- 07jul05
fiber tests show that 174Khz ripple comes from the fiber
receiver.
- 06jul05
fractional occurrence of rfi for jun05
- 01jul05
sbn receiver warms up during sband radar run.
- 25jun05
td 12 releases after lightning strike.
- 23jun05
lbw resonance polA at 1400 Mhz
- 21jun05
lbw cals measured 01jun05 after polB cal cable fixed.
- 12jun05
1428Mhz birdie in alfa 18may05 to 23may05.
- 10jun05
alfa beammaps from spider scans (on 3c138 taken jan05)
- 09jun05
time for sband dewar to cool down.
- 07jun05
16 khz comb in the 327Mhz receiver.
- 05jun05
wapp fits file usage.. updated to look at all 3 discs..
- 05jun05
updated a2010 (alfalfa) monitor data for month of may05.
- 03jun05
time for lbw dewar to cool down.
- 02jun05
fractional occurrence of rfi for may05
- 31may05
weekly report of problems started.. 25may05
- 31may05
galfa pixel4a,6a were switched (before 02may05 until 31may05)
- 30may05
distomat 4 loses power 6pm. ac filter at input to canister burns
up...
- 28may05
lbw stokes calibration. included date on 4 more sources
jul04-apr05.
- 26may05
alfa pixel4a still has 173Khz ripple with amp of .09K.
- 20may05
monitor disc usage by project in the wapp spectral line disc
directory.
- 19may05
discrete gain jumps and a1.2 hz crosshead oscillation in the lbw
gain.
- 09may05
comparing lbw, alfa performance at OH with the 10db amps in.
- 07may05
added p2030 user project page.
- 07may05
p2030 pulsar data feb05 thru apr05 shows aerostat compression
with new amps.
- 05may05
modeling the periods created by the aerostat radar.
- 05may05
put aerostat radar info in one page
- 02may05
fractional occurrence of rfi for apr05.
- 02may05
updated system performance for apr05.
- 19apr05
drifting of the wapp clocks (ntp off)
- 19apr05
ntpd time jumps on the wapps when pulsar search p2030 run.
- 12apr05
a2010 1 second total power vs position in strip for apr05.
- 11apr05
p2030 pulsar data shows compression with new rf amps and wapps
not starting together.
- 08apr05
a2010 drift scan data (apr05) with new amp in pix2 show radar
compression.
- 06apr05
a2010 monitoring. add data for apr05.
- 06apr05
10db gain amp placed in front of alfa pixel2a,b mixer.
- 05apr05
fractional occurrence of rfi for mar05
- 04apr05
Gain stability of the alfa system using a2010 drift scan data.
- 04apr05
updated system performance for mar05
- 02apr05
Comparing the system performace of alfa and lbw at oh
frequencies.
- 30mar05
faa 1350 radar harmonics appear in alfa surveys.
- 29mar05
data taken 18nov04 show 1400Mhz birdies in distomats 3,4.
- 24mar05
1430 Mhz birdie in alfa pixel0b.
- 17mar05
430 linefeed gain curve.
- 17mar05
430 linefeed calibration page added.
- 17mar05
pointing model 16 for the carriage house made and installed.
- 14mar05
include 430 ch in system performance plots.
- 13mar05
updated platform temperature measurements from 2000 through
feb05.
- 10mar05
TsysA/TsysB ratio for alfa.. added plots vs ra,dec,l, and b.
- 08mar05
TsysA/TsysB ratio for alfa spectral wapp data. 01nov04 thru
28feb05.
- 08mar05
description of computing tsys from the tsys archive.
- 05mar05
fractional occurrence of rfi for feb05 updated.
- 04mar05
updated system performance for feb05
- 04mar05
alfa pwr jumps (a2010,a1963), slow recovery from faa
radar. Problem not in downstairs iflo.
- 01mar05
327 receiver gain curves for data taken mar03 to feb05.
- 01mar05
327 system performance for data used to fit the gain curves.
- 27feb05
26feb05: punta salinas radar does not stay in mode A
- 18feb040
lband multibeam system. Show rotation rate vs parallactic angle
and maximum integration times.
- 6feb05
how much does the az/za move if we drift in J2000 coordinates.
- 16feb05
hpbw vs lamda for all receivers. data taken 2004.
- 15feb05
fractional occurrence of rfi oct,nov,dec04,jan05. Cleaned
up alfa plots. Thanks to mmd.
- 13feb05
hi frequency system performance after reflectors adjusted.
- 04feb05
dome cam limit switch replaced with switch/strikers. New limits
measured.
- 03feb05
td12 moves erratically and then goes into foldback.
- 02feb05
updated system performance thru 31jan05
- 02feb05
fractional occurrence of rfi for jan05
- 31jan05
Single pol/board wapp spectral line observing only provides 1
pol.
- 24jan05
For tsys monitoring, included plots for TsysPolA - TsysPolB
- 23jan05
recomputed sysperformance for alfa using the new cal values.
- 23jan05
plots of alfa Tcal vs frequency. Data taken 30sep04.
- 21jan05
look at cb,cbh,xb gain before and after reflector adjustments.
- 18jan05
distomat failures, include plots az vs za,and az vs hour of day.
- 18jan05
The 60Mhz harmonic birdie put to rest...
- 11jan05
430 mapping project a1731. Variation of Tsys.
- 11jan05
430 Mhz TsysPolA-TsyspolB drifting for 2003, 2004.
- 06jan05
updated 29nov04 alfa rotation angle to fix pointing error in
computation. new angle 1.55 deg.
- 05jan05
updated system performance for dec04. (now includes alfa).
2004 (top)
- 14dec04
drifting 60 Mhz birdie seen at 60 Mhz in the dome with a
spectrum analyzer
- 12dec04
fractional occurrence of rfi for nov04.
- 10dec04
updated system performance plots for nov04.
- 12dev04
Calibration scans with alfa, rotation angle offset, and beammaps
(a first pass).
- 01dec04
Look at how the fits header locations are computed.
- 23nov04
9186Mhz birdie at xband coming from oscillation in sbw dewar.
- 22nov04
proj a1552.. cordrift. Negative ghost of galaxy in strip caused
by bandpass correction.
- 20nov04
19nov04 data shows 60Mhz birdie is coming from the dewar
monitoring system.
- 20nov04
18nov04 data: 60Mhz harmonics jump ever 2.67 seconds (1 ms data
with wapps).
- 19nov04
12nov04 data: 60 Mhz birdies are coming thru
- 18nov04
1367/1382 radar birdies were punta salinas with chn 18 left on.
- 16nov04
tiedown 4 trouble with power supply and then motor overspeed.
- 12nov04
azimuth dependence of faa radar using alfa and 64 usec dumps.
- 08nov04
fractional occurrence of rfi for oct04.
- 07nov04
28oct04 data shows lband drifting birdies seen at lband
and 430 Mhz.
- 07nov04
system performance plots for oct04.
- 06nov04
FCC rules for broad band over power lines access (oct04)
- 06nov04
FCC reallocation of the 1390,1430,1670,2385 Mhz bands (dec01)
- 03nov04
galfa spectrometer time stamp stability.
- 01nov04
spider scan (cima) positions updated so all strips pass thru
center of source.
- 31oct04
alfa precursor a1956. driftchase pattern position errors from
21oct04.
- 20oct04
alfa precursor a1956. Test data taken 28sep04. Positions start
of scan off.
- 17oct04
sbn 1.2 hz oscillation in polB. same frequency as the dewar
refrig.
- 09oct04
430 Mhz filter in front of dewar removes the tv intermods.
- 09oct04
fractional occurrence of rfi for sept04
- 09oct04
updated system performance for sep04
- 07oct04
updated alfa jumping from radar. Pulsar data shows it is
the faa radar.
- 07oct04
punta salinas, faa, and aerostat radars total power vs time
using alfa.
- 29sep04
cband polA oscillations and Tsys jumping around.
- 27sep04
radar is causing the total power in some alfa beams to increase.
- 24sep04
checking the ra,dec J2000 coordinates in the header after the
16sep04 changes.
- 22sep04
mathew's 430 test feed performance after lowering it 3 feet.
- 21sep04
Phase delay needed in the line feed for focusing of the ao dish.
- 17sep04
wind speed and platform oscillation frequencies for tropical
storm jeanne.
- 14sep04
a1946 alfa drift scans. Jumps in the total power 06sep04.
- 13sep04
cross scans on 3c286 using Jmathews test feed on the carriage
house.
- 12sep04
cal values for lband wide measured 06aug04, installed 12sep04.
- 11sep04
added sbw to plot off gain vs freq,za for all receivers.
- 10sep04
installed sbw gain curve for data taken 15may04 - 01sep04.
- 10sep04
created sbw calibration page
- 10sep04
a description of the gain curves used at ao.
- 08sep04
1380 Mhz Birdie in pixel 0b of alfa data.
- 08sep04
174 khz ripple in alfa pixel4a data as seen by wapps and galfa.
- 07sep04
carl heiles' notes on the galfa spectrometer.
- 02sep04
proj
a1946,alfa
month
of
aug.
Check
that
brds
have
consistent
hdr
info
and
that
time
steps
are
1 sec between records.
- 01sep04
updated sysperformance for aug04
- 01sep04
Tsys alfa using cal on/offs from aug04.
- 27aug04
the stability of the async cal mode of cima.
- 25aug04
updated the alfa pixel numbering drawing to be correct.
- 22aug04
a1946 drift scan thru a source. data timestamp ok, rotation
angle ok (this page updated 27aug04).
- 22aug04
a1946 cross scans after beam order changed.
- 20aug04
a1946 alfa precursor proposal. data taken 19aug04 (cross scans).
- 16aug04
calibration scans with alfa at 1665 Mhz.
- 13aug04
dome vibrations while moving at slew rate at low za.
- 10aug04
measuring the alfa calvalues at 1665 Mhz using sky and absorber.
- 09aug04
fractional occurrence of rfi for jul04.
- 08aug04
updated system performance for jul04.
- 02aug04
ripples in lbw near 1390 and 1400 Mhz (a1861:may04,a1914:aug04).
- 01aug04
faa radar compression of the alfa system.
- 20jul04
updated sbw system performance to new cal diode values.
- 18jul04
measured sbw cal values using sky and absorber.
- 13jul04
fractional occurrence of rfi for jun04
- 08jul04
Using mean,median, and robust averge to remove Tsys in
cormapping routines.
- 04jul04
updated system performance thru jun04.
- 03jul04
updated gain curves all receivers with new sbh fit
- 03jul04
sbh gain curves from feb04 thru jun04 data.
- 03jul04
sbh system performance used to compute gain curves feb04-jun04
- 02jul04
1400 Mhz birdie coming from 100Mhz distribution/wapps in control
room building.
- 01jul04
azimuth swing looking at 1400 Mhz birdies with high spectral
resolution.
- 30jun04
added high freq resolution dynamic spectra of 1374 Mhz birdie.
- 29jun04
distomat birdies (30may04) still present after communications
changed from copper to fiber.
- 28jun04
lbw polB cal jumps on 23apr04
- 22jun04
drifting birdies in lb system with 60 Mhz spacing.
- 21jun04
summary of 430 Mhz birdies may/jun04 compiled by chris salter.
- 17jun04
xband performance above 10Ghz.
- 17jun04
extend xband cal value vs freq plot to 12.3 ghz.
- 10jun04
added fractional occurrence of rfi for may04
- 05jun04
pointing error as we lose tension in the tiedowns.
- 04jun04
added system performance for may04
- 29may04
added fractional occurrence rfi for apr04
- 28may04
1374.495 Mhz birdie (a1861).
- 28may04
1400 Mhz birdie (a1861).
- 27may04
started list of monitor data that is backed up.
- 23may04
Compare az,za model offsets with horn survey.
- 23may04
the az,za model offsets for the "other" receivers relative to
sbn from model15.
- 18may04
aas plots for jagadheep
- 15may04
pointing model 15
- 13may04
updated tiedown 4 limit switch positions after jack replaced.
- 09may04
results of the 08may04 horn survey.
- 08may04
added blurb on turret encoder.
- 04may04
Alfa Tsys measured using the old iflo system.
- 04may04
alfa cal values measured in the lab.
- 03may04
horn survey(02may04) after moving the horns.
- 01may04
surveying the tertiary (01may04)
- 29apr04
horn survey results 28apr04.
- 27apr04
resonance in alfa receiver at 1537.7 Mhz.
- 26apr04
Alfa Tsys/Tcal using the old iflo system.
- 12apr04
computing alfa wedge position on the floor.
- 09apr04
quick look lbw performance after kevlar cable tensioning.
- 07apr04
rcvr system performance thru 01apr04
- 07apr04
fractional occurrence of rfi for feb04
- 27mar04
xband gain curves mar03 thru sep03
- 18mar04
xband tsys vs frequency.
- 18mar04
installed new cal values for xband receiver.
- 14mar04
new cal values for sband high..
- 14mar04
added description on controlling the gregorian cals...
- 14mar04
moved cal values from home page to a page of its own.
- 10mar04
sbh gain curve fit using data taken 15may03 thru 30sep03
- 10mar04
system performance sbh for gain curve dataset.
- 09mar04
sbn gain curve fit using data taken 01mar03 thru 29feb04
- 09mar04
system performance sbn for gain curve dataset.
- 08mar04
cband gain curves using data from may03 thru feb04.
- 08mar04
system performance cband may03 thru feb04. data used for gain
fit.
- 08mar04
rcvr system performance thru 29feb04 (cband hi now
included).
- 08mar04
fractional occurrence of rfi for feb04
- 04mar04
looking at the alfa rf/if processor using a noise source.
- 29feb04
fit gain curve to cbh feb04 calibration data.
- 27feb04
notes to myself.. computing Trcvr.
- 25feb04
distomat failure vs az,za
- 25feb04
platform corner height vs za, torque. Will ALFA be a problem??
- 23feb04
looking at 6.7 ghz masers with the cband high receiver.
- 23feb04
added cbh to tsys monitoring output.
- 18feb04
updated sysperformance page thru 01feb04.
- 18feb04
corrected parallactic angle picture.
- 17feb04
tsys vs frequency for cband hi.
- 17feb04
a first look at the rfi in the cband hi receiver band.
- 16feb04
(on-off)/off on 3C454.3 covering the entire cbhi band.
- 16feb04
a first look at cband high calibration.
- 15feb04
update the horn survey page to include the 11feb04 and 14feb04
horn theodolite surveys.
- 12feb04
atod , signal conditioning board description for turret.
- 01feb04
fractional occurrence of rfi for jan04
- 25jan04
computed gain curve for xband oct05 to 25jan03.
- 25jan04
moving the 327 receiver on the turret floor.
- 25jan04
moving the xband receiver on the turret floor.
- 24jan04
moving the 610 Mhz rcvr to a new turret position.
- 06jan04
1422.5 Mhz birdie from channel 54 transmitter.
- 05jan04
defining the sign of the pointing error for turret and calib
scans.
- 02jan04
updated system performance thru 31dec03
- 02jan04
fractional occurrence of rfi for dec03
2003 (top)
- 15dec03
updated system performance thru 15dec03.
- 15dec03
fractional occurrence of rfi for nov03.
- 12dec03
sband transmitter page. Include TTI module i/o usage.
- 04dec03
372Khz comb in 327 rcvr from sbTx console in dome.
- 02dec03
372Khz comb in 327 rcvr coming from inside the SbandTx room.
- 25nov03
372Khz comb in 327 rcvr.
- 12nov03
fractional occurrence of rfi for oct03.
- 12nov03
weight distribution on the turret rotary floor.
- 09nov03
updated system performance for oct03.
- 02nov03
updated gain curve page to include new lbw curve.
- 02nov03
lbw gain curves computed using data from mar03 thru sep03.
- 30oct03
1417.5 birdie from color camera 5 in the dome.
- 30oct03
added az,gr,ch limits to little know facts
- 21oct03
Reflector survey 2 (oct01) added normal errors along north/south
direction.
- 19oct03
Reflector survey 2 (oct01) results.
- 15oct03
azimuth encoder jump 14oct03.
- 14oct03
1289.7 Mhz radar (punta salinas??)
- 12oct03
time stability of the iridium signal.
- 10oct03
sband wide cal values from 15aug03 data.
- 05oct03
updated system performance for month of sept03
- 05oct03
add fractional occurrence of rfi for sept.
- 25sep03
theodolite survey results from jul03.
- 24sep03
first holography outputs from aips (via slava).
- 17sep03
xband: resonances in receiver.
- 14sep03
xband: compare sefd new and old horn
- 14sep03
Compare xband TsysB/TsysB using cals and source.
- 14sep03
Tsys vs frquency for xband before, after new horn.
- 10sep03
added fractional occurrence of rfi for august
- 10sep03
increasing the isolation for iflo monitor port 1a.
- 10sep03
added tsys ascii output description on tsys mon page.
- 08sep03
updated system performance for month of august.
- 02sep03
added routine arch_getmap() to input a cormap from the data
archive.
- 01sep03
holography,plots of phase vs time,za , correction at the peaks.
- 23aug03
a1659 sband hi CH ons only. Example of processing this type of
data.
- 12aug03
bs116b vlba holography run. telescope offsets for each pattern.
- 07aug03
system performance. reran data 01mar03 to aug03 with new flux
fits.
- 05aug03
added fractional occurrence of rfi for july.
- 31jul03
430 dome birdies with 20 Mhz filter installed.
- 28jul03
added tiedown page and failure log.
- 21jul03
xb pointing offsets from calibration runs
- 21jul03
new xb cal values installed (sky/absorber measurements)
- 17jul03
ri glitches in digitizer Q1.
- 13jul03
compare kildal and german cortes horns on lbw.
- 02jul03
added fractional occurrence of rfi for jun03
- 02jul03
updated system performance page thru 30jun03
- 01jul03
alfa IF tested 15may03 using lbw receiver.
- 01jul03
az swings looking at polarized Tsys with sbh.
- 27jun03
start collecting rfi birdies in a list by freq.
- 27jun03
collect various rfi az spins together for comparison.
- 27jun03
az spin with 430 Mhz birdies. tv station intermods.
- 25jun03
recomputed lbn gain curves for sep01 thru 20aug02.
- 04jun03
fractional occurrence of rfi for may03.
- 03jun03
xband. polb variations for a1704 recomb lines.
- 01jun03
system performance updated thru 31may03. lbn,lbw,sbw,sbh,cb,xb.
- 29may03
added correlator drifts for data thru 27may03
- 19may03
lbw stokes (new lbw) src:B1040+123,B1228+307,B2121+248
- 18may03
cband cal values measured 16may03 installed.
- 18may03
cband cal values 22apr03 to 04may03 installed.
- 17may03
updated links for "using idl to process spectral line data"
- 15may03
resonances in the sbh receiver.
- 14may03
sband high cals recalibrated.
- 09may03
added corsavbysrc for idl automated on/off processing.
- 08apr03
.3 kelvin polarized tsys vs az,za using lbw.
- 07apr03
fractional occurrence of rfi for apr03
- 06may03
lbw 1290,1350 radars. 1ms detect. show rotation drift and
compression.
- 04may03
system performance 430 dome.
- 04may03
system performance lbw. x102 runs 02may03
- 04may03
system performance 327 rcvr.
- 03may03
system performance 610 rcvr.
- 02may03
system performance. sbn mar03 to may03
- 02may03
system performance. xband apr02 to feb03.
- 02may03
system performance. lbw mar03 to may03
- 01may03
added system performance page. Started with lbn,lbw,sbn,cb,xb.
- 29apr03
IFLO compression by the FAA 1350/1330 radars using lbw,lbn.
- 28apr03
dewar temperatures. year to date is now being updated for new
data.
- 27apr03
lbn system performance 01sep01 thru 20aug02 (with updated
calvalues).
- 23apr03
poor isolation in the downstairs iflo buffer amplifier monitor
ports.
- 22apr03
sefd vs for lbw using 3C138
- 17apr03
resonances in the new lbw receiver.
- 13apr03
az,zenith angle dependence of 1350,1290, and aerostat radars.
- 09mar03
fractional occurrence of rfi for mar03
- 07apr03
slow drift in gain of sbc1 polB in interim correlator.
- 06apr03
radar power in the downstairs iflo.
- 31mar03
creating harmonics in an amplifier.
- 26mar03
tv station power from hilltop monitoring. 1998 to current date.
- 20mar03
galfa talk on rfi.
- 15mar03
pitch,roll model from 28feb03 tilt sensor run (after dome move).
- 15mar03
tilt sensor results from 28feb03 run (after dome move)
uncalibrated..
- 14mar03
jumps in interim correlator sbc 2 power levels.
- 08mar03
fractional occurrence of rfi feb03
- 07mar03
finished model14. included variogram, verifying model, other
rcvr offsets.
- 07mar03
added model offsets each rcvr to turret page.
- 05mar03
pointing model14 results.
- 04mar03
turret room temps for feb03
- 02mar03
Compare focus curves before after dome move.
- 02mar03
focus curve 28feb03 after the dome move.
- 27feb03
lbw gain curves 21nov02 to 17feb03 after old lbw phase center
was moved.
- 27feb03
lbw system performance after the old lbw phase center was
changed.
- 27feb03
1371,1387.2 radar..
- 22feb03
compare focus curves taken after 01dec02.
- 21feb03
focus curves sbn: CTA21, 3C138 taken 18feb03 after shimming.
- 20feb03
make a pitch,roll model from the 19feb03 tilt sensor swings.
- 20feb03
added tilt sensor run from 19feb03
- 20feb03
theodolite results from feb03 survey
- 20feb03
tilt sensor results from feb03 survey
- 16feb03
ao9 survey of dome 12feb03 after first shimming cycle.
- 15feb03
tilt sensors compare pitch roll 05feb03 (before shimming), and
12feb03 after 1st shimming.
- 14feb03
Azswings, za strips tilt sensor data 04aug01 during aug01 a09
survey.
- 12feb03
Azswings, zastrips tilt sensor data 05feb03 before elevation
rail shimming.
- 07feb03
fractional occurrence of rfi jan03
- 06feb03
compare pitch,roll, focus 09aug01 with 04feb03 (before
shimming).
- 06feb03
turret room temps for jan03
- 02feb03
tilt sensor angle with the turret floor.
- 02feb03
tilt
sensor
measurements
of
the 1,3az terms (16dec02).
- 31jan03
yr2002 cable tensions vs platform
hght,temp,tdPositions. data and fits.
- 11jan03
added 10jan03 focus curve to comparison of most recent focus
curves.
- 11jan03
added cband focus curve taken 10jan03.
- 05jan03
fractional occurrence of rfi dec02
- 01jan03
turret room temps for dec02.
2002 (top)
- 30dec02
sbn narrow focus curve done 16dec02.
- 30dec02
sbn sefd ratio shows that tsysA,B difference is real.
- 23dec02
max sky az,za sky velocity vs za.
- 18dec02
added graphic monitor of 16K stage dewar temperatures.
- 15dec02
lb,sb,cb gain.. showing sources used to compute gain curves.
- 11dec02
Updated Alfa postAmp page showing gain needed in postamps.
- 06dec02
plot of gain curves with new cband data.
- 06dec02
cband gain curve after horn focus.data 28aug02 thru 30nov02.
- 06dec02
cband system performance 28aug02(after horn focus) thru 30nov02.
- 06dec02
fractional occurrence of rfi for nov02
- 05dec02
compare gain/sefd sep00 to aug02..lband,sband,cband.
- 04dec02
turret room temperatures for nov02.
- 04dec02
100 Mhz frequency reference unlocks (25nov02).
- 02dec02
lbw system performance change after horn moved.
- 24nov02
lbw horn focus change.
- 20nov02
added idl routines for ao pointing.
- 07nov02
Using the calOn - calOff to remove the bandpass shape.
- 05nov02
cross scans on seti@home feed taken 01jul02.
- 04oct02
primary reflector errors measure oct01 (east half from lynn
bakers plots)
- 04nov02
turret room temperatures for oct02.
- 04nov02
fractional occurrence of rfi for oct02.
- 28oct02
add rfi hilltop monitor page. started record of high freq
receiver.
- 23oct02
410 Mhz birdies from the navy transmitter in aguada.
- 20oct02
added monitoring of distomat failures by month.
- 13oct02
stokes reduction. reran lbw,lbn,sbw sources.
- 07oct02
added more programs to the telescope monitoring page.
- 05oct02
spectrum analyz plots of lbw dewar output taken 18sep02.
- 05oct02
updated lbn stokes data to 01oct02.
- 03oct02
some notes on the multibeam power levels for the post amp
chassis.
- 03oct02
fractional occurrence of rfi for sep02
- 03oct02
turret room temperatures for sep02.
- 29sep02
gain curves for various receivers.
- 29sep02
Rms noise versus integration time using the interim correlator
and a noise src.
- 28sep02
spike in central channel of correlator bandpasses.
- 25sep02
correlator bandpasses using a noise source.
- 25sep02
Added to: comparing measured and computed bandpass shapes...
- 16sep02
azimuth encoder jumps.
- 10sep02
lbn gain curves after horn focus change (data 21aug02 thru
09sep02)
- 10sep02
lbn system performance 21aug02 thru 09sep02 (after horn focus
change).
- 06sep02
Comparing the measured and computed bandpass shapes of the
digital filters.
- 03sep02
turret room temperatures for aug02.
- 03sep02
fractional occurrence of rfi for aug02.
- 30aug02
corbl: idl routine to interactivly baseline a correlator
dataset.
- 29aug02
cband gain curves from data taken sep01 to aug02
- 28aug02
gain before/after moving the cb horn so phase center is at
focus.
- 27aug02
Telescope monitor displays running from multicast memory.
- 24aug02
Tsys vs turret position before, after lbn focus change.
- 23aug02
added corcalib idl routine. Calibrate on only scan with
calonoff.
- 22aug02
moving the lbn horn to the correct phase center.
- 20aug02
added ao idl correlator routines to the downloadable software
page.
- 19aug02
added selectrcv program documentation. Creates a subset of a
datafile by rcvr.
- 19aug02
added srcTimes program documentation. compute lst rise,set
times.
- 18aug02
added bToJ B1950->J2000 program documentation.
- 17aug02
added handy filter programs documentation. so far just cordmp.
- 16aug02
system performance for cband sep01 through aug01.
- 11aug02
327 calibration runs
- 09aug02
327 receiver breakdown of tsys contributions.
- 05aug02
fractional occurrence of rfi for jul02.
- 05aug02
added turret room temperatures for jul02
- 04aug02
updated tertiary connection point locations from 020204 survey.
- 01aug02
dome pitch,roll while td's track the average platform height.
- 01aug02
resonances in the cband receiver.
- 31jul02
resonances in the sbh receiver.
- 24jul02
2
MHz
comb
in
327 Mhz receiver 327 2mhz comb from 430 waveguide power monitor.
Gasket
material
on
huffman
boxes
degrade
with
time.
- 30jul02
added page to download software. First routine is cordmp.
- 27jul02
2 hz gain variation in xband from the rcvr crosshead.
- 25jul02
2 mhz comb in 327 receiver.
- 24jul02
xband drift scans.
- 10jul02
added idl routines to look at agc (az,gr,ch) data.
- 08jul02
pointing errors with and without the dome hydraulic brake.
- 07jul02
fractional occurrence of rfi for jun02.
- 04jul02
turret room temperatures for jun02
- 28jun02
gregorian dome hydraulic brake performance
- 25jun02
torques needed to move the dome 2<->20 degrees za.
- 22jun02
dome amplifier glitching when changing directions.
- 14jun02
TsysA-TsysB variation with azimut using Cband receiver at low
za.
- 13jun02
horn spacing for multi-beam system.
- 11jun02
Measure loss from 19cm and 25 cm horn offset using lbn (for
multi beam tests).
- 08jun02
platform and bowl temperatures 2000-2002.
- 06jun02
faa radar harmonics when 1350Mhz radar not within digital filter
bands.
- 04jun02
turret room temperatures for may02.
- 04jun02
fractional occurrence of rfi for may02.
- 03jun02
compared lbn cal values 28jan02 computed by phil and karen and
compare sefd lbw,lbn.
- 29may02
brake valve on dome clogs while testing.. data 21may02
- 27may02
color camera biridies at lband (taken 24may02) with camera #1
not installed.
- 15may02
lbn sefd over entire receiver band.
- 15may02
moved lbw calibration list to separate page.
- 13may02
corrected focus curves for kildal phase center error.
- 11may02
lbw cal value. Rms polB worse than polA (probably the ripple in
polB)
- 10may02
lbw polB ripple. switched dewar output cables. ripple in dewar.
- 10may02
fractional occurrence of rfi for apr02
- 09may02
turret room temperatures for month apr02 added.
- 01may02
Ripple in lband wide pol B still present.
- 28apr02
distomat birdie at 1399.82 reappears.
- 26apr02
added lbn to "best" turret position by tracking sources.
- 25apr02
updated stokes data for cband sep01->15apr02
- 24apr02
update stokes data for sbw sep01->>15apr02
- 23apr02
updated stokes data for sbn jun01->15apr02
- 23apr02
updated stokes data for lbw jan01->15apr02
- 22apr02
Look for "best" turret position by tracking sources.
- 20apr02
updated stokes data for lbn thru apr02
- 19apr02
added corstokes idl routine to intensity calibrate stokes data.
- 15apr02
added mmgetarchive idl routine to extract calibration data.
- 15apr02
added sband high calibration runs for 08apr02,11apr02.
- 14apr02
tsys versus turret position. tsys bumps.. updated for sbh.
- 09apr02
rfi from color cameras in the dome reappears.
- 07apr02
resonances in the lband wide omt.
- 06apr02
added system temperature plots: daily, yeartoday, by year.
- 05apr02
added fractional occurrence of rfi for mar02
- 04apr02
added turret room temperatures for march. New Ac now on..
- 03apr02
added horn survey results.
- 03apr02
include sband high focus curve in comparison of focus curves.
- 03apr02
added sband high focus curve.
- 26mar02
added dewar temperature monitoring.
- 26mar02
added monitoring info section it's easier to find some of this
data.
- 24mar02
comparison of gain before and after reflector adjusted.
- 23mar02
sband high (3-4 ghz) Tsys and calvalues versus frequency.
- 21mar02
added idl routines to display the dewar temperatures.
- 19mar02
focusing with the tertiary.
- 17mar02
sbh cal values measured at the rcvr test shack.
- 11mar02
added lbw focus curve.
- 11mar02
added lbn focus curve.
- 11mar02
comparing focus curves. added lbn,lbw, sbn from mar02.
- 11mar02
redid sband narrow focus curve to check that tertiary hadn't
moved (it didn't)
- 09mar02
fractional occurrence of rfi for feb02. xband now included.
- 08mar02
added feb02 turret room temperatures.
- 08mar02
sbw focus curve.
- 08mar02
comparing focus curves. Added sbw, separate plot of most recent
curves.
- 07mar02
hc3n,hc7n in tmc1 using xband receiver.
- 05mar02
comparing focus curves
- 04mar02
lbn cals. added values measured 28jan02 (but not installed as
cal values)
- 04mar02
lbn cals. added values measured 06dec01 (but not installed as
cal values).
- 03mar02
added az swing to the 1400 Mhz birdie.
- 02mar02
Computed "other" cband cals using ratio to hi correlated cal.
- 28feb02
1400 Mhz birdie (not the distomats).
- 28feb02
HC3N lines at 9097,9098, and 9100 Mhz possible detection at
xband.
- 24feb02
22feb02 Tsys vs turret position lbn,lbw,sbn,sbw,cb,xb. Shows
Tsys bump.
- 22feb02
21feb02: xband calibration run 3C138,3C286, and B1040+123.
- 21feb02
xband:10 beammaps (turret scan) J0237+288 after pointing offsets
added.
- 20feb02
history of turret positions and pointing az,za offsets by model.
- 20feb02
how many points does it take to move pointing model to a (non
sbn) receiver.
- 19feb02
cband pitch,roll correlation with coma.
- 18feb02
xband calib run on J2253+161,3C48 on 02jan02
- 18feb02
added cor idl plotting routines mmplotgtsb,mmplotcsme.
- 17feb02
xband pnterr, tsys,sefd,beammaps on J2253+161,J0137+331 (3C48)
using turret scans.
- 15feb02
show the fraction of time the remy radar spent at 1270,1290 Mhz
for 2001.
- 12feb02
measured the radar parameters for the 1270 radar at remy.
- 10feb02
added fractional occurrence of rfi for
327,430.610,lband,sband,cband through jan02.
- 09feb02
added to the 1az term discussion for the horizontal offsets of
dome.
- 07feb02
xband focus curve after shimming receiver.
- 06feb02
added history of turret positions for receivers.
- 05feb02
compute shims needed for pitch,roll, and focus correction.
- 04feb02
added turret room temps for dec01,jan02
- 04feb02
simulate tiedown motion after shimming the rails.
- 31jan02
cband focus curve taken 23jan02
- 30jan02
sband narrow focus curve taken 20jan02
- 30jan02
little known facts.
- 30jan02
tiedown motion needed to remove pitch,roll,focus error.
- 27jan02
pitch,roll of telescope using theodolite and tilt sensors (with
2d fit)
- 23jan02
added fractional occurrence of rfi for dec01.
- 17jan02
Use survey azimuth, encoder az differences to compute platform
offset from AO9
- 12jan02
first shot at adding an azimuth encoder table to model 13.
- 09jan02
pointing model 13 (jan02).
- 04jan02
prelimary plots of ao9 survey 09aug01.
- 03jan02
multifreqency measurements of 3C38,3C138 sep01->dec01
- 02jan02
xband beammaps on J2253+161 using turret scans.
- 02jan02
xband pointing error and beamwidths from turret scans.
- 01jan02
xband focus curve after receiver installed (24dec01).
2001 (top)
- 26dec01
large ripple in 430 Mhz gregorian dome receiver.
- 23dec01
cband sefd vs za preview from 22dec01 calibration run.
- 23dec01
xband installation. added rfi plots. First out of focus light:
gain,sefd,beammaps.
- 21dec01
xband installation. added power levels in if1 and rfi seen at
xband.
- 20dec01
xband installation. turret strips for turret position,
calvalues, tsys versus frequency.
- 13dec01
Updated measuring cal with sky/absorber to include reflection
coefficient.
- 10dec01
added software modification history page.
- 07dec01
added fractional occurrence of rfi for month of nov01
- 03dec01
sband narrow calibration sep-nov01
- 03dec01
lband wide calibration sep-nov01
- 21nov01
stokes: added parameter definitions, srclist, updated cband to
19nov01.
- 19nov01
lbw gain/tsys calibration from Sep01.
- 11nov01
added idl documentation for generic routines.
- 08nov01
1381 gps L3 birdie seems to be on all the time.
- 03nov01
Added turret room temp for month of october.
- 31oct01
Timer module in air conditioner compressor causing rfi.
- 12oct01
sband wide cal values measured 20sep01.
- 10oct01
Description of measuring cals on the telescope using absorber
and sky.
- 10oct01
1350 Mhz rfi tone when blanker in use.
- 09oct01
added images to illustrate radar harmonics at 260 IF.
- 08oct01
resonances in the sband wide receiver (20sep01 data)
- 06oct01
New airconditioner, cycle louvres, blower, compressor (5oct01
data).
- 05oct01
lbn image with magnetic relay installed on new ac compressor.
- 04oct01
Radar harmonics created in the 2nd IF
- 02oct01
idl routines to look at fractional rfi occurrence interactively.
- 02oct01
monthly update. fractional rfi occurrence for sept.
- 02oct01
monthly update. turret room temperature for September.
- 24sep01
lbn stokes. Redid with no mueller matrix correction (previous
had used it).
- 22sep01
lbn image showing jumps when aircondioner cycled
on/off (from 20sep01).
- 22sep01
rfi color camera on service platform. before/after shielding.
- 21sep01
pointing strips on sbw. updated za model offsets.
- 21sep01
spectrum of the sband dewar output.
- 19sep01
fractional rfi occurrence by freq. added aug01 and new
yeartodat. (lband).
- 17sep01
fractional rfi occurrence by frequency plots added. (lband).
- 15sep01
lbn total power jumps coming from new airconditioner compressor
running.
- 14sep01
added turret room temp by hour and month (for 2001)
- 13sep01
updated info on a1459 baseline ripples.
- 10sep01
1% baseline ripple in polB lbwide that's been there for months.
- 09sep01
a1459 baseline ripple. lbn 1350-1420. 3min on/offs
- 29aug01
azimuth servo failures. motor torques build up while stopped
(24aug01).
- 15aug01
rfi from color camera on the service platform (13aug01).
- 15aug01
rfi from tertiary motors and electronics (02/08aug01).
- 14aug01
cband gain/sefd after panels in northern half of dish adjusted
(02aug01).
- 13aug01
pointing and gain changes from tertiary motor install (sbn).
- 02aug01
rfi from tertiary cpu control box (mounted in the dome).
- 02aug01
tertiary motors installed on tertiary. encoder tracking error
for a sine wave.
- 28jul01
Added cabling page under hardware. Shows fiber connections.
- 14jul01
software documentation to input and plot platform height/temp
from distomats.
- 10jul01
tertiary simulator output.
- 03jul01
Problems with using the lbw hybrid to make circular
polarization.
- 01jul01
cband gain/sefd with pitch,roll,focus correction after SE quad
adjusted.
- 01jul01
3C48 gain at cband after SE quad of dish adjusted. Compare with
02feb01.
- 01jul01
Fractional gain versus surface rms by freq. ruze curves.
- 30jun01
full stokes reduction of 3C48 after se quad of dish adjusted.
- 30jun01
Added system modification history page.
- 21jun01
carriage house hydraulic brake. Measure free fall velocity (no
motor).
- 19jun01
added jan-jun01 lbn,cband,sbn stokes calibration data.
- 15jun01
stokes polarization calibration.
- 14jun01
lband narrow filter bank calibration.
- 12jun01
temperature dependence lbw cals. calon/off tracking blank sky.
- 08jun01
Slow rise time for cband cal to go from 95 to 100% of maximum.
- 08jun01
Carriage moving uphill/downhill with hydraulic brake on 1 motor.
- 02jun01
Interim correlator. Adjusting power levels with large out of
band signal.
- 28may01
pointing error and gain before and after tertiary platform
installed.
- 21may01
azimuth encoder rack gear error (measured 21jan00).
- 12may01
azimuth blown by wind while brakes on (06may01).
- 11may01
finished additions to rfi 1407 birdie
- 11may01
added tv station information page.
- 02may01
lband wide power levels at output of the dewar. also update the
1407 birdie info.
- 28apr01
1407 rfi birdie from tv station.
- 24apr01
idl routines to look at rfi from hilltop mo>
- 16apr01
compare turret scans with heiles calibration on CTA21
- 11apr01
the multi meter in the new rfi box.
- 06apr01
platform motion as we pull down equally on all 3 tiedowns.
- 27mar01
Added pointing page. started with model 12 preliminary results.
- 23feb01
pitch,roll, focus correction at multiple frequencies on 3C48,
CTA21(feb01)
- 21feb01
cband calibration after surface adjusted jan01.
- 21feb01 moved sband calibration after surface
adjustment from 11. misc. to sband calibration.
- 16feb01 updated focus gain loss. Routine was
using wrong fwhm. switched to 2 sigma.
Mainly
changed
what
gain
would
look
like
after
we
did
the
prf
correction.
- 14feb01
add cband focus curve after jan01 reflector adjustment
- 14feb01
add sband focus curve after jan01 reflector adjustment
- 14feb01
added focus curve page
- 08feb01
plot measured surface errors primary versus azimuth,za.
- 31jan01
calibrate slide rule potentiometer versus encoder for tertiary.
- 28jan01
cband cals measured using absorber and sky
- 27jan01
430 dome looking for rfi from vittara jeep parked next to the
control room.
- 27jan01
lband rfi at 1300 mhz from the distomats.
- 20jan01
added rfi removal techniques with frank briggs paper..
- 17jan01
x111 image display of rfi data.
- 17jan01
tertiary ramp dac -10 to 10 counts. plot dac val, velocity and
force.
- 17jan01
Added tertiary entries. tracking error, force calibration,
friction.
- 09jan01
Telescope gain after the eastern half of the reflector has been
adjusted.
2000 (top)
- 29dec00
Added lbn cal value measurements.
- 28dec00
tsys
during az swings with dome at 19 deg za. Shows 120 deg tsys
bumps and sun getting into the beam.
- 27dec00
rfi measurement. lbn 1420Mhz, azswing dome at 19 degrees, 3mhz
bw.
- 26dec00
Added rfi measurements page.
- 01dec00
System temperature versus turret angle offset.
- 27oct00
rotary floor room temperature versus hour of day.
- 10oct00
source tracks across the dish by declination and hour angle.
- 10oct00
added lband multi beam link to home page and paralactic rotation
rate vs dec.
- 05oct00
327 calibration. turret strips showing turret position, and
polA,polB offsets
- 02oct00
lband fit gain(az,za). Changed to color. included fit to only za
dependence.
- 23sep00
sband narrow calibration. Sefd, gain, tsys, and fit to (az,za).
-
added
coef's to cband calibration.
-
converted
cta21 multifreq measurements to color
-
converted
lband (az,za) plots to color.
- 21sep00
cband calibration. Sefd, gain, tsys and fits to (az,za).
- 20sep00
lband fit gain(az,za). Changed to average gain.Updated
plots,text
- 04sep00
CTA21 multifreq measurements gain. compute surface error via
ruze formula.
- 28aug00
gain stability vs temp (after sband xmit experiment).
- 25aug00
610 receiver calibration
- 23aug00
lband fit gain(az,za) and comparison with pitch,roll,focus
errrors
- 17aug00
rfi: radar ipps memo from 1997/1999
- 17aug00 started logging modification dates
<\ul>